### v4

# Universal Shuffle Asymptotics: Sharp Privacy Analysis in the Gaussian Regime

Abstract
We develop a sharp, experiment-level privacy theory for amplification by shuffling in the Gaussian regime: a fixed finite-output local randomizer with full support and neighboring binary datasets differing in one user. The goal is structural: to isolate the fixed full-support regime in which Gaussian/GDP is the right limit object and to make the controlling constants explicit. We first prove exact likelihood-ratio identities for shuffled histograms and a complete conditional-expectation linearization theorem with explicit typical-set remainders. We then derive sharp Jensen–Shannon divergence expansions, identifying the universal leading constant $I_{\pi}/(8n)$ for proportional compositions and emphasizing the correct fixed-composition covariance $\Sigma_{\pi}=(1-\pi)\Sigma_{0}+\pi\Sigma_{1}$. Next we establish equivalence to Gaussian Differential Privacy with Berry–Esseen bounds and obtain the full limiting $(\varepsilon,\delta)$ privacy curve. Finally, for unbundled multi-message shuffling we give an exact degree-$m$ likelihood ratio, asymptotic GDP formulas, and a strict bundled-versus-unbundled comparison. Further results include Local Asymptotic Normality with quantitative Le Cam equivalence, exact finite-$n$ privacy curves, a boundary Berry–Esseen theorem for randomized response, and a frequency-estimation application.
[…]
MSC 2020: 62G10, 94A17, 68P27.
Keywords: differential privacy, shuffle model, privacy amplification, Gaussian differential privacy, Jensen–Shannon divergence, local asymptotic normality.
1 Introduction
The shuffle model of differential privacy interpolates between the local and central models by inserting a trusted shuffler between users and an analyst. Each of $n$ users holds a bit $x_{i}\in\{0,1\}$, applies a local randomizer $W:\{0,1\}\to\Delta(\mathcal{Y})$, and releases a message in a finite alphabet $\mathcal{Y}=\{1,\dots,d\}$. The neighboring input model is binary, but the output alphabet $\mathcal{Y}$ is an arbitrary fixed finite set. The shuffler outputs a uniformly random permutation of the $n$ messages, so the analyst observes only the permuted multiset, equivalently the histogram $N\in\mathbb{Z}_{\geq 0}^{d}$ of message counts. For a dataset with exactly $k$ ones, denote the induced histogram law by $T_{n,k}$. By exchangeability, $T_{n,k}$ depends only on $k$, and any neighboring pair of datasets induces some neighboring pair $(T_{n,k},T_{n,k+1})$.
A large literature provides amplification-by-shuffling bounds, and for some concrete mechanisms finite-$n$ guarantees can also be computed numerically. Those methods answer a certification question for fixed parameters: what privacy guarantee do we get for this mechanism at this $n$? They do not identify the governing regime, tell us when a Gaussian/GDP approximation is actually the right one, or explain which features of the local channel control privacy level and convergence speed. For protocol design, mechanism comparison, and sharpness questions, that structural information matters as much as the certificate itself. This paper gives a self-contained treatment of the Gaussian regime: fixed finite-output local randomizers with full support, where neighboring shuffle experiments are asymptotically equivalent (in Le Cam distance) to a one-dimensional Gaussian shift. Two structural inputs drive the analysis: (1) exact likelihood-ratio identities that reduce shuffled comparisons to controlled functionals of a score statistic; and (2) a complete conditional-expectation linearization theorem with explicit typical-set remainders, proved via a lattice Edgeworth ratio lemma.
An important distinction (fixed-composition covariance).
For proportional compositions ($k/n\to\pi$), the histogram is a fixed-composition sum: exactly nkn-k users sample from $W_{0}$ and $k$ users sample from $W_{1}$. The correct per-user covariance is therefore
$\Sigma_{\pi}=(1-\pi)\Sigma_{0}+\pi\Sigma_{1},\qquad\Sigma_{b}=\mathrm{diag}(W_{b})-W_{b}W_{b}^{\top},$
not the i.i.d. mixture covariance $\mathrm{diag}(f_{\pi})-f_{\pi}f_{\pi}^{\top}$ of $f_{\pi}=(1-\pi)W_{0}+\pi W_{1}$. The difference is the rank-one term $\pi(1-\pi)vv^{\top}$ (with $v=W_{1}-W_{0}$), which inflates the mixture covariance and leads to a smaller (overly optimistic) Fisher constant. All proportional-composition first-order constants (JSD, GDP/LAN parameters, and privacy curves) are stated here in terms of the correct $\Sigma_{\pi}$.
Companion paper (Part II).
When local privacy parameters grow with $n$ (e.g. randomized response with $\varepsilon_{0}\asymp\log n$), Lindeberg conditions can fail and Gaussian limits can break. The companion paper (Part II) identifies non-Gaussian critical and super-critical limits (Poisson/Skellam/point-process regimes) and provides explicit privacy curves in those regimes.
Related work (brief).
Differential privacy was introduced by Dwork et al. [8]; see Dwork–Roth [9] for a textbook treatment. Amplification by shuffling was initiated and developed in Erlingsson et al. [11] and Cheu et al. [6], with many subsequent refinements. Feldman–McMillan–Talwar [13] provide a simple nearly optimal amplification guarantee, and Balle–Bell–Gascón–Nissim [3] introduce the privacy blanket method. Numerical inversion of shuffle-amplification bounds is also used in practice; the ENPA white paper [1] is one concrete example. Rényi divergence accounting was developed by Mironov [20] and extended to subsampled settings by Balle–Barthe–Gaboardi [2] and Wang–Balle–Kasiviswanathan [28]. Concentrated differential privacy [10, 5], composition theorems [16, 24], and numerical privacy accounting via FFT and characteristic functions [18, 30] provide complementary tools. These upper-bound and numerical methods address certification for given mechanisms. Our goal here is different: to identify the fixed full-support regime itself, the correct Gaussian/GDP limit in that regime, and the constants that govern approximation and comparison. The exact linear likelihood ratio identity for the canonical shuffle pair appears (in equivalent form) in Girgis et al. [14] as part of their shuffle RDP analysis. Independent and concurrent work by Takagi–Liew [25] introduces a “shuffle index” that corresponds to $\chi^{2}(W_{1}\|W_{0})$ for the canonical pair; a preliminary version of the present work appeared on Zenodo on January 3, 2026 (DOI: 10.5281/zenodo.18112202). Our framework extends to general compositions, unbundled multi-message models, and exact privacy curves.
Road map.
Section 2 sets up the model, notation, and the corrected fixed-composition covariance. Section 3 records exact likelihood-ratio identities. Section 4 proves conditional-expectation linearization with quantitative typical-set remainders. Section 5 derives sharp Jensen–Shannon expansions. Section 6 establishes Gaussian Differential Privacy equivalence and the limiting $(\varepsilon,\delta)$ curve. Section 7 develops Local Asymptotic Normality and quantitative Le Cam equivalence. Section 8 gives exact finite-$n$ privacy-curve identities and computable bounds. Section 9 treats unbundled multi-message shuffling and bundled-versus-unbundled comparison. Section 10 analyzes the Gaussian boundary for randomized response, and Section 11 presents a frequency-estimation application. Section 12 concludes with open problems.
2 Model, definitions, standing assumptions
2.1 Shuffle mechanism and histogram experiments
Fix a finite output alphabet $\mathcal{Y}=\{1,\dots,d\}$. A (single-message) local randomizer is a channel
$W:\{0,1\}\to\Delta(\mathcal{Y}),\qquad x\mapsto W_{x}(\cdot),$
where each $W_{x}$ is a probability mass function on $\mathcal{Y}$. Given a dataset $D=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$, users release independent messages $Y_{i}\sim W_{x_{i}}$, and the shuffler reveals only the histogram
$N=(N_{y})_{y\in\mathcal{Y}}\in\mathbb{Z}_{\geq 0}^{d},\qquad N_{y}=\sum_{i=1}^{n}\mathbf{1}\{Y_{i}=y\},\qquad\sum_{y\in\mathcal{Y}}N_{y}=n.$
For $k\in\{0,\dots,n\}$ let $T_{n,k}$ be the law of $N$ when exactly $k$ users have input $1$ and nkn-k have input 0. Neighboring datasets (Hamming distance $1$) correspond, by symmetry, to neighboring compositions $(T_{n,k},T_{n,k+1})$.
Remark 2.1 (Reduction to composition-indexed experiments).
By exchangeability, any neighboring pair of binary datasets induces $(T_{n,k},T_{n,k+1})$ for some $k$, justifying the focus on composition-indexed experiments.
2.2 Likelihood ratios, privacy loss, divergences, and privacy curves
For distributions $P,Q$ on a common finite space with $Q\ll P$, let $L=dQ/dP$ be the likelihood ratio and $\Lambda=\log L$ the privacy loss. We use the Jensen–Shannon divergence
$\mathrm{JSD}(P\|Q):=\tfrac{1}{2}\mathrm{KL}(P\|M)+\tfrac{1}{2}\mathrm{KL}(Q\|M),\qquad M:=\tfrac{1}{2}(P+Q).$
We also use the chi-square divergence $\chi^{2}(Q\|P):=\mathbb{E}_{P}[(L-1)^{2}]$ and the Rényi divergence [20]
$D_{\alpha}(Q\|P):=\frac{1}{\alpha-1}\log\mathbb{E}_{P}[L^{\alpha}],\qquad\alpha>1.$
For $(\varepsilon,\delta)$-type statements we work with the one-sided privacy curve
$\delta_{Q\|P}(\varepsilon):=\sup_{A}\{Q(A)-e^{\varepsilon}P(A)\}=\mathbb{E}_{P}[(L-e^{\varepsilon})_{+}],\qquad\varepsilon\geq 0,$
and the two-sided curve $\delta(\varepsilon):=\max\{\delta_{Q\|P}(\varepsilon),\delta_{P\|Q}(\varepsilon)\}$.
We also use trade-off functions for binary experiments (Section 7).
2.3 Standing support assumptions
We separate minimal absolute-continuity assumptions from the stronger full-support condition used for bounded-privacy-loss arguments.
Definition 2.2 (Minimum mass parameters).
Define
$\delta_{\star}:=\min_{y\in\mathcal{Y}}W_{0}(y),\qquad\delta_{\mathrm{full}}:=\min_{y\in\mathcal{Y}}\min\{W_{0}(y),W_{1}(y)\}.$
Remark 2.3.
If $\delta_{\star}>0$, then $T_{n,1}\ll T_{n,0}$ and likelihood ratios are finite. For results controlling $\Lambda=\log L$ (Berry–Esseen, GDP, LAN), we assume $\delta_{\mathrm{full}}>0$, which makes $\Lambda$ bounded. If some symbols have zero probability under one input, the analysis can be run after restricting $\mathcal{Y}$ to the common support $\{y:W_{0}(y)>0,\ W_{1}(y)>0\}$, provided that the singular component (symbols with $W_{0}(y)>0=W_{1}(y)$ or vice versa) is accounted for separately: such events contribute infinite privacy loss in the corresponding direction and must be included when computing $(\varepsilon,\delta)$ curves. The Gaussian regime treated in this paper is confined to the absolutely continuous (full-support) case.
Remark 2.4 (Guide to assumptions).
Results requiring only $\delta_{\star}>0$ include the likelihood-ratio identities (Section 3), the sharp canonical JSD expansion (Theorem 5.4), and exact finite-$n$ privacy curves (Section 8). Results requiring $\delta_{\mathrm{full}}>0$ include linearization (Theorem 4.1), the proportional-composition JSD constant (Theorem 5.5), Berry–Esseen and GDP equivalence (Section 6), LAN (Section 7), and unbundled multi-message analysis (Section 9).
2.4 Correct covariance for fixed compositions and Fisher constant
Write $W_{0},W_{1}$ as vectors in $\mathbb{R}^{d}$, define the difference
$v:=W_{1}-W_{0}\in\mathbb{R}^{d},\qquad\mathbf{1}^{\top}v=0,$
and the within-type categorical covariances
$\Sigma_{b}:=\mathrm{diag}(W_{b})-W_{b}W_{b}^{\top},\qquad b\in\{0,1\}.$
For proportional compositions $k/n\to\pi\in[0,1]$, the shuffle histogram is a sum of independent one-hot vectors with fixed composition (exactly nkn-k draws from $W_{0}$ and $k$ draws from $W_{1}$). The correct per-user covariance in this fixed-composition experiment is therefore
$\Sigma_{\pi}:=(1-\pi)\Sigma_{0}+\pi\Sigma_{1}.$ (1)
This is the quantity that governs CLT variances and Fisher information in the proportional-composition Gaussian regime.
It is sometimes convenient to compare (1) to the multinomial-mixture covariance of the i.i.d. mixture $f_{\pi}=(1-\pi)W_{0}+\pi W_{1}$:
$\Sigma^{\mathrm{mix}}_{\pi}:=\mathrm{diag}(f_{\pi})-f_{\pi}f_{\pi}^{\top}.$
A direct computation gives the rank-one correction
$\Sigma^{\mathrm{mix}}_{\pi}=\Sigma_{\pi}+\pi(1-\pi)\,vv^{\top}.$ (2)
In particular, $\Sigma^{\mathrm{mix}}_{\pi}$ is larger than $\Sigma_{\pi}$ in the PSD order, so using $\Sigma^{\mathrm{mix}}_{\pi}$ leads to an optimistically small Fisher constant.
Let $T:=\{x\in\mathbb{R}^{d}:\mathbf{1}^{\top}x=0\}$ be the tangent space, and let $\Sigma_{\pi}^{+}$ denote the Moore–Penrose pseudoinverse (inverse on $T$ and 0 on $\mathrm{span}\{\mathbf{1}\}$). Define the Fisher constant
$I_{\pi}:=v^{\top}\Sigma_{\pi}^{+}v.$ (3)
When $v=0$ (i.e. $W_{0}=W_{1}$), $I_{\pi}=0$ for all $\pi$ and privacy is perfect; we therefore assume $v\neq 0$ throughout.
Remark 2.5 (Canonical case: $I_{0}=\chi^{2}$).
For the canonical pair ($\pi=0$), $\Sigma_{0}=\mathrm{diag}(W_{0})-W_{0}W_{0}^{\top}$ and $\Sigma_{0}^{+}$ acts on $T$ as $\mathrm{diag}(1/W_{0})$. Since $v\in T$, we have $I_{0}=v^{\top}\Sigma_{0}^{+}v=\sum_{y}v(y)^{2}/W_{0}(y)=\chi^{2}(W_{1}\|W_{0})$, the chi-square divergence. This connects the Fisher constant of Sections 57 with the chi-square appearing in the canonical results of Sections 35.
Proposition 2.6 (Relation to multinomial Fisher proxy).
Let $\pi\in(0,1)$ and write $f_{\pi}=(1-\pi)W_{0}+\pi W_{1}$. Define the multinomial Fisher proxy
$I_{f}:=\sum_{y\in\mathcal{Y}}\frac{v(y)^{2}}{f_{\pi}(y)}.$
Then
$I_{\pi}=\frac{I_{f}}{1-\pi(1-\pi)I_{f}},\qquad\text{so in particular}\qquad I_{\pi}\geq I_{f},$ (4)
with equality if and only if $\pi\in\{0,1\}$ or $v=0$ (i.e. $W_{0}=W_{1}$).
Proof.
Let $\Sigma^{\mathrm{mix}}_{\pi}=\mathrm{diag}(f_{\pi})-f_{\pi}f_{\pi}^{\top}$. On $T=\{x:\mathbf{1}^{\top}x=0\}$ this matrix is invertible, and for any $u\in T$,
$u^{\top}(\Sigma^{\mathrm{mix}}_{\pi})^{+}u=u^{\top}\mathrm{diag}(1/f_{\pi})\,u=\sum_{y\in\mathcal{Y}}\frac{u(y)^{2}}{f_{\pi}(y)},$
since the rank-one ambiguity in the pseudoinverse acts only on $\mathrm{span}\{\mathbf{1}\}$ and $\mathbf{1}^{\top}u=0$. In particular, $I_{f}=v^{\top}(\Sigma^{\mathrm{mix}}_{\pi})^{+}v$. Using (2), on $T$ we have $\Sigma_{\pi}=\Sigma^{\mathrm{mix}}_{\pi}-\pi(1-\pi)vv^{\top}$. Apply the Sherman–Morrison formula on $T$ with $A=\Sigma^{\mathrm{mix}}_{\pi}|_{T}$ and $u=\sqrt{\pi(1-\pi)}\,v$:
$(\Sigma_{\pi}|_{T})^{-1}=A^{-1}+\frac{A^{-1}uu^{\top}A^{-1}}{1-u^{\top}A^{-1}u}.$
Taking the quadratic form in direction $v$ gives $I_{\pi}=v^{\top}\Sigma_{\pi}^{+}v=I_{f}/(1-\pi(1-\pi)I_{f})$. Since $\Sigma_{\pi}|_{T}\succ 0$, the denominator lies in $(0,1]$, hence $I_{\pi}\geq I_{f}$ with equality if and only if $\pi(1-\pi)I_{f}=0$, i.e. $\pi\in\{0,1\}$ or $v=0$. ∎
2.5 Notation table
Symbol Meaning
$n$ number of users
$k$ number of ones in the dataset; composition $\pi=k/n$
$\mathcal{Y}=\{1,\dots,d\}$ finite output alphabet
$W_{0},W_{1}$ local randomizer output laws on inputs $0,1$
$N=(N_{y})$ released histogram
$T_{n,k}$ histogram law for composition $k$
$L_{n,k},\Lambda_{n,k}$ likelihood ratio $dT_{n,k+1}/dT_{n,k}$ and privacy loss $\log L_{n,k}$
$w(y),r(y)$ pointwise LR and increment: $w=W_{1}/W_{0}$, r=w1r=w-1
$v$ difference vector $v=W_{1}-W_{0}$
$\Sigma_{0},\Sigma_{1}$ within-type covariances $\mathrm{diag}(W_{b})-W_{b}W_{b}^{\top}$
$\Sigma_{\pi}$ fixed-composition covariance $(1-\pi)\Sigma_{0}+\pi\Sigma_{1}$
$\Sigma_{\pi}^{+}$ Moore–Penrose pseudoinverse on $T=\{x:\mathbf{1}^{\top}x=0\}$
$I_{\pi}$ Fisher constant $v^{\top}\Sigma_{\pi}^{+}v$
$\mu_{n}(\pi)$ GDP/LAN parameter $\sqrt{I_{\pi}/n}$; for $k=0$: $\mu_{n}=\sqrt{\chi^{2}/n}$ (Remark 2.5)
$\mu_{n,\pi}(m)$ unbundled GDP parameter $\sqrt{mI_{\pi}/n}$ (Section 9)
$a_{n}$ RR boundary scaling $e^{\varepsilon_{0}}/n$
$q_{n}$ RR flip probability $q_{n}=(1+e^{\varepsilon_{0}})^{-1}$
$D(t)$ JSD pointwise functional: $\frac{1}{2}\log\frac{2}{1+t}+\frac{t}{2}\log\frac{2t}{1+t}$
$\delta_{Q\|P}(\varepsilon)$ one-sided privacy curve $\mathbb{E}_{P}[(L-e^{\varepsilon})_{+}]$
$\mathcal{T}_{P,Q}(\alpha)$ trade-off function of the binary experiment $(P,Q)$
Table 1: Frequently used notation.
3 Exact likelihood ratio identities
3.1 Canonical neighboring pair: linear likelihood ratio
We first treat the canonical neighboring datasets: $k=0$ (all zeros) versus $k=1$ (one one).
Definition 3.1 (Ratio increment and score average).
Assume $\delta_{\star}>0$ and define
$w(y):=\frac{W_{1}(y)}{W_{0}(y)},\qquad r(y):=w(y)-1.$
Under $T_{n,0}$ the messages $Y_{1},\dots,Y_{n}$ are i.i.d. with law $W_{0}$. Let $Z_{i}:=r(Y_{i})$ and
$U_{n}:=\frac{1}{n}\sum_{i=1}^{n}Z_{i}.$
Lemma 3.2 (Exact shuffle likelihood ratio for the canonical pair).
Assume $\delta_{\star}>0$. For every histogram $N$,
$L_{n,0}(N):=\frac{dT_{n,1}}{dT_{n,0}}(N)=\frac{1}{n}\sum_{y\in\mathcal{Y}}N_{y}\frac{W_{1}(y)}{W_{0}(y)}=1+U_{n}.$
Moreover $0\leq L_{n,0}\leq\delta_{\star}^{-1}$ and hence $U_{n}\in[-1,\delta_{\star}^{-1}-1]$ almost surely under $T_{n,0}$.
Proof.
Fix $N$ with $\sum_{y}N_{y}=n$. Under $T_{n,0}$ the histogram pmf is multinomial:
$T_{n,0}(N)=\frac{n!}{\prod_{y}N_{y}!}\prod_{y\in\mathcal{Y}}W_{0}(y)^{N_{y}}.$
Under $T_{n,1}$, condition on the output $Y_{\star}=y$ of the changed user; the remaining n1n-1 messages have histogram $N-e_{y}$ and are distributed as under $T_{n-1,0}$. Thus
$T_{n,1}(N)=\sum_{y\in\mathcal{Y}}W_{1}(y)\,\frac{(n-1)!}{(N_{y}-1)!\prod_{z\neq y}N_{z}!}\,W_{0}(y)^{N_{y}-1}\!\!\prod_{z\neq y}W_{0}(z)^{N_{z}}.$
Dividing by $T_{n,0}(N)$ yields $\sum_{y}(N_{y}/n)\,W_{1}(y)/W_{0}(y)$, which equals $1+U_{n}$ after rewriting $N_{y}=\sum_{i}\mathbf{1}\{Y_{i}=y\}$. The bounds follow from $W_{1}(y)\leq 1$ and $W_{0}(y)\geq\delta_{\star}$. ∎
3.2 General neighboring pair: conditional-expectation likelihood ratio
For general $k$, the likelihood ratio is typically nonlinear in the histogram.
Lemma 3.3 (General-$k$ likelihood ratio as a conditional expectation).
Fix $k\in\{0,\dots,n-1\}$. Let $P=T_{n,k}$ and $Q=T_{n,k+1}$. Let $Y_{\star}$ be the message of the unique user whose bit changes between the datasets, so $Y_{\star}\sim W_{0}$ under $P$ and $Y_{\star}\sim W_{1}$ under $Q$. Then, for every histogram $N$,
$L_{n,k}(N):=\frac{dQ}{dP}(N)=\mathbb{E}\!\left[\left.\frac{W_{1}(Y_{\star})}{W_{0}(Y_{\star})}\ \right|\ N\right]=1+\mathbb{E}[r(Y_{\star})\mid N].$
Proof.
Condition on $Y_{\star}=y$. Under either hypothesis, the remaining n1n-1 users consist of $k$ ones and nk1n-k-1 zeros, hence their shuffled histogram has the same law, say $T_{n-1,k}$. Therefore
$P(N)=\sum_{y}W_{0}(y)\,T_{n-1,k}(N-e_{y}),\qquad Q(N)=\sum_{y}W_{1}(y)\,T_{n-1,k}(N-e_{y}).$
Dividing yields
$\frac{Q(N)}{P(N)}=\sum_{y}\frac{W_{1}(y)}{W_{0}(y)}\cdot\frac{W_{0}(y)\,T_{n-1,k}(N-e_{y})}{P(N)}=\mathbb{E}\!\left[\left.\frac{W_{1}(Y_{\star})}{W_{0}(Y_{\star})}\ \right|\ N\right].$
4 Complete conditional-expectation linearization
Lemma 3.3 reduces the proportional-composition analysis to controlling the posterior expectation $\mathbb{E}[r(Y_{\star})\mid N]$ uniformly on a typical set. The main theorem below provides the required linearization with explicit remainder bounds.
4.1 Statement
Throughout this section, assume full support $\delta_{\mathrm{full}}>0$ and fixed alphabet size $d$.
For $k/n\to\pi\in(0,1)$, in the decomposition $N=X_{1}+S$ with $S=\sum_{i=2}^{n}X_{i}$, the relevant finite-$n$ proportion for the (n1)(n-1)-sum is
$\pi_{n}:=\frac{k}{n-1}.$
Define $f_{\pi_{n}}=(1-\pi_{n})W_{0}+\pi_{n}W_{1}$ and the fixed-composition covariance
$\Sigma_{\pi_{n}}:=(1-\pi_{n})\Sigma_{0}+\pi_{n}\Sigma_{1}.$
Let $s_{n}:=\Sigma_{\pi_{n}}^{+}v\in T$.
Theorem 4.1 (Conditional expectation linearization).
Assume full support $\delta_{\mathrm{full}}>0$ and fixed alphabet size $d$. Let $\pi_{n}=k/(n-1)$ and define $s_{n}=\Sigma_{\pi_{n}}^{+}v$. Define the typical set
$A_{n}:=\left\{\|N-\mathbb{E}[N]\|_{\infty}\leq n^{5/8}\right\}.$
There exists $c=c(d,\delta_{\mathrm{full}})>0$ such that $\mathbb{P}(A_{n}^{c})\leq 2d\exp(-cn^{1/4})$. Moreover, on $A_{n}$ the conditional expectation admits the linearization
$\mathbb{E}[r(Y_{\star})\mid N]=\frac{1}{n}s_{n}^{\top}\bigl(N-\mathbb{E}[N]\bigr)+R_{n}(N),$
where $\sup_{N\in A_{n}}|R_{n}(N)|\leq Cn^{-3/4}$ for a constant $C=C(d,\delta_{\mathrm{full}})$.
4.2 Proof of Theorem 4.1
Proof.
We follow a four-step argument: (1) a typical-set concentration; (2) a local CLT ratio lemma for the remaining sum; (3) Bayes’ rule and a first-order posterior expansion; (4) computation of the conditional expectation and identification of the Fisher direction.
Step 1: typical set and reduction to d1d-1 coordinates.
Let $\tilde{x}\in\mathbb{R}^{d-1}$ denote the vector obtained by dropping the last coordinate of $x\in\mathbb{R}^{d}$, and write $\tilde{N},\tilde{S},\tilde{X}_{i}$ accordingly. Then $\tilde{N}=\tilde{X}_{1}+\tilde{S}$ with $\tilde{X}_{1}=\tilde{e}_{Y_{\star}}$ where $\tilde{e}_{y}=e_{y}\in\mathbb{R}^{d-1}$ for $y\leq d-1$ and $\tilde{e}_{d}=0$. Each $N_{y}$ is a sum of independent Bernoulli variables with success probabilities in $\delta_{\mathrm{full}},1-\delta_{\mathrm{full}}]$, so Bernstein’s inequality yields
$\mathbb{P}(|N_{y}-\mathbb{E}[N_{y}]|>n^{5/8})\leq 2\exp(-c_{y}n^{1/4})$
with $c_{y}=c_{y}(\delta_{\mathrm{full}})>0$. A union bound over $y\in\mathcal{Y}$ gives the stated tail bound with $c=\min_{y}c_{y}$.
Step 2: local CLT ratio lemma for $\tilde{S}$.
Let $\tilde{\mu}_{n}:=\mathbb{E}[\tilde{S}$]. On $A_{n}$, we have $\|\tilde{N}-\tilde{\mu}_{n}\|_{\infty}\leq n^{5/8}$ (up to an $O(1)$ discrepancy from dropping one coordinate). The key local ratio estimate is Lemma A.2 in Appendix A, applied to the lattice-valued sum $\tilde{S}$ (a sum of n1n-1 independent one-hot vectors with uniform full support). Note the correspondence with Lemma A.2: a shift $u=-\tilde{e}_{y}$ in $\mathbb{R}^{d-1}$ corresponds to $t=e_{d}-e_{y}\in\mathbb{R}^{d}$ with $\mathbf{1}^{\top}t=0$ and $\|t\|_{1}=2$, so the zero-sum constraint of Lemma A.2 is automatically satisfied. (When $y=d$, the shift is $\tilde{e}_{d}=0$ and the ratio is trivially $1$.) For any fixed $u\in\mathbb{Z}^{d-1}$ with $\|u\|_{\infty}\leq d$ and any $\tilde{s}$ with $\|\tilde{s}-\tilde{\mu}_{n}\|_{\infty}\leq n^{5/8}$,
$\log\frac{\mathbb{P}(\tilde{S}=\tilde{s}+u)}{\mathbb{P}(\tilde{S}=\tilde{s})}=-\frac{1}{n-1}\,u^{\top}\tilde{\Sigma}_{\pi_{n}}^{-1}(\tilde{s}-\tilde{\mu}_{n})-\frac{1}{2(n-1)}u^{\top}\tilde{\Sigma}_{\pi_{n}}^{-1}u+\rho_{n}(\tilde{s},u),$ (5)
where $\tilde{\Sigma}_{\pi_{n}}$ is the $(d-1)\times(d-1)$ covariance minor of $\Sigma_{\pi_{n}}$ and
$\sup_{\|\tilde{s}-\tilde{\mu}_{n}\|_{\infty}\leq n^{5/8}}\sup_{\|u\|_{\infty}\leq d}|\rho_{n}(\tilde{s},u)|\leq C_{0}n^{-3/4}.$ (6)
Step 3: Bayes formula for the posterior of $Y_{\star}$ and first-order expansion.
Fix $\tilde{n}$ with $\tilde{N}=\tilde{n}$. Bayes’ formula and $\tilde{N}=\tilde{X}_{1}+\tilde{S}$ give, for $y\in\mathcal{Y}$,
$\mathbb{P}(Y_{\star}=y\mid\tilde{N}=\tilde{n})=\frac{W_{0}(y)\,\mathbb{P}(\tilde{S}=\tilde{n}-\tilde{e}_{y})}{\sum_{y^{\prime}\in\mathcal{Y}}W_{0}(y^{\prime})\,\mathbb{P}(\tilde{S}=\tilde{n}-\tilde{e}_{y^{\prime}})}.$ (7)
On $A_{n}$, apply (5) with $\tilde{s}=\tilde{n}$ and $u=-\tilde{e}_{y}$ and exponentiate. Define the tilt parameter
$\theta_{n}:=\frac{1}{n-1}\tilde{\Sigma}_{\pi_{n}}^{-1}(\tilde{N}-\tilde{\mu}_{n})\in\mathbb{R}^{d-1}.$
Then (5) implies
$\mathbb{P}(\tilde{S}=\tilde{n}-\tilde{e}_{y})=\mathbb{P}(\tilde{S}=\tilde{n})\exp\Bigl(\tilde{e}_{y}^{\top}\theta_{n}+\eta_{n,y}(N)\Bigr),$
where
$\eta_{n,y}(N):=-\frac{1}{2(n-1)}\tilde{e}_{y}^{\top}\tilde{\Sigma}_{\pi_{n}}^{-1}\tilde{e}_{y}+\rho_{n}(\tilde{n},-\tilde{e}_{y}).$
[…]
$\mathbb{P}(Y_{\star}=y\mid N)=\frac{W_{0}(y)\exp(\tilde{e}_{y}^{\top}\theta_{n}+\eta_{n,y}(N))}{Z_{n}(\theta_{n})},\qquad Z_{n}(\theta_{n}):=\sum_{y^{\prime}\in\mathcal{Y}}W_{0}(y^{\prime})\exp(\tilde{e}_{y^{\prime}}^{\top}\theta_{n}+\eta_{n,y^{\prime}}(N)).$
Taylor expanding the exponential and the normalizer up to first order in $\theta_{n}$ and using $\eta_{n,y}=O(n^{-3/4})$ uniformly yields
$\mathbb{P}(Y_{\star}=y\mid N)=W_{0}(y)\Bigl(1+(\tilde{e}_{y}-\tilde{W}_{0})^{\top}\theta_{n}\Bigr)+O(n^{-3/4}),$ (8)
uniformly over $y\in\mathcal{Y}$ and $N\in A_{n}$, where $\tilde{W}_{0}\in\mathbb{R}^{d-1}$ denotes the first d1d-1 coordinates of $W_{0}$.
Step 4: compute $\mathbb{E}[r(Y_{\star})\mid N]$ and identify $s_{n}=\Sigma_{\pi_{n}}^{+}v$.
Using (8) and boundedness of $r$ under full support,
$\mathbb{E}[r(Y_{\star})\mid N]=\sum_{y\in\mathcal{Y}}r(y)\,\mathbb{P}(Y_{\star}=y\mid N)=\sum_{y\in\mathcal{Y}}r(y)W_{0}(y)\Bigl(1+(\tilde{e}_{y}-\tilde{W}_{0})^{\top}\theta_{n}\Bigr)+O(n^{-3/4}).$
The constant term vanishes because $\sum_{y}r(y)W_{0}(y)=\sum_{y}(W_{1}(y)-W_{0}(y))=0$. For the linear term, note that $r(y)W_{0}(y)=W_{1}(y)-W_{0}(y)=v(y)$, hence $\sum_{y}r(y)W_{0}(y)\tilde{e}_{y}=\tilde{v}$ and the $\tilde{W}_{0}$ term cancels by the same zero-sum. Therefore,
$\mathbb{E}[r(Y_{\star})\mid N]=\tilde{v}^{\top}\theta_{n}+O(n^{-3/4})=\frac{1}{n-1}\tilde{v}^{\top}\tilde{\Sigma}_{\pi_{n}}^{-1}(\tilde{N}-\tilde{\mu}_{n})+O(n^{-3/4}).$
It remains to rewrite the right-hand side in $d$ dimensions. Since $\tilde{\mu}_{n}=\mathbb{E}[\tilde{S}]=\mathbb{E}[\tilde{N}]-\tilde{W}_{0}$, we have $\tilde{N}-\tilde{\mu}_{n}=(\tilde{N}-\mathbb{E}[\tilde{N}])+\tilde{W}_{0}$. The (d1)(d-1)-dimensional inverse $\tilde{\Sigma}_{\pi_{n}}^{-1}$ corresponds to the pseudoinverse $\Sigma_{\pi_{n}}^{+}$ on $T$ after eliminating one coordinate (see (36) in Appendix A), so
$\tilde{v}^{\top}\tilde{\Sigma}_{\pi_{n}}^{-1}(\tilde{N}-\tilde{\mu}_{n})=v^{\top}\Sigma_{\pi_{n}}^{+}(N-\mathbb{E}[N])+\Delta_{0},$
where $\Delta_{0}:=\tilde{v}^{\top}\tilde{\Sigma}_{\pi_{n}}^{-1}\tilde{W}_{0}$ is a constant depending only on $W_{0},W_{1},\pi_{n}$, and $d$ (not on $N$). Dividing by n1n-1 gives an $O(n^{-1})$ contribution, which is negligible compared to $n^{-3/4}$. Finally, replacing $(n-1)^{-1}$ by $n^{-1}$ in the leading term produces an error of order $(n(n-1))^{-1}\cdot O_{P}(\sqrt{n})=O_{P}(n^{-3/2})$ on $A_{n}$, again negligible. This yields the stated expansion with remainder bounded by $Cn^{-3/4}$ on $A_{n}$. ∎
4.3 Refined remainder on a smaller typical set
The $O(n^{-3/4})$ remainder in Theorem 4.1 is dictated by the window $\|N-\mathbb{E}[N]\|_{\infty}\leq n^{5/8}$, which corresponds to standardized deviations of order $n^{1/8}$ in the local Edgeworth polynomial. On a smaller window, one can refine the remainder nearly to $n^{-1}$ (up to polylog factors).
Proposition 4.2 (Refined remainder on a $\sqrt{n\log n}$ window).
Under the assumptions of Theorem 4.1, define the smaller typical set
$B_{n}:=\left\{\|N-\mathbb{E}[N]\|_{\infty}\leq\sqrt{n\log n}\right\}.$
Then there exist constants $c,C>0$ (depending only on $d$ and $\delta_{\mathrm{full}}$) such that
$\mathbb{P}(B_{n}^{c})\leq 2d\,n^{-c},$
and on $B_{n}$,
$\mathbb{E}[r(Y_{\star})\mid N]=\frac{1}{n}s_{n}^{\top}\bigl(N-\mathbb{E}[N]\bigr)+\widetilde{R}_{n}(N),\qquad\sup_{N\in B_{n}}|\widetilde{R}_{n}(N)|\leq C\frac{\log^{2}n}{n}.$
Proof.
Repeat the proof of Theorem 4.1, replacing Bernstein’s inequality on scale $n^{5/8}$ by the scale $\sqrt{n\log n}$ (yielding $\mathbb{P}(B_{n}^{c})\leq 2dn^{-c}$), and replacing the ratio lemma remainder by its small-window specialization in Appendix A. On $B_{n}$, the standardized deviation $z=(s-\mu)/\sqrt{n}$ is $O(\sqrt{\log n})$ rather than $O(n^{1/8})$, and the Edgeworth polynomial gradient bounds improve the ratio remainder from $O(n^{-3/4})$ to $O((\log n)/n)$. Propagating this through the posterior tilt and Taylor expansion yields the stated $O(\log^{2}n/n)$ bound. ∎
Remark 4.3 (Connection to sharp $1/n$ remainder conjectures).
Proposition 4.2 shows that near-$n^{-1}$ remainders (up to $\log^{2}n$) hold on a window whose complement has only polynomially small probability. Achieving a clean uniform $O(n^{-1})$ remainder globally requires (i) controlling Edgeworth polynomial derivatives up to standardized deviations of order $n^{1/8}$ and (ii) careful accounting of rare-event contributions outside the typical set.
4.4 Uniform sharpness for $k=\Theta(n)$
The next theorem resolves the open problem stated above: one can obtain a uniform $O(n^{-2})$ JSD remainder for all proportional compositions $k=\Theta(n)$ without needing a pointwise $O(n^{-1})$ linearization remainder on an exponentially likely window. The key is to (i) use an exact linear-regression identity that pins down the $1/n$ projection coefficient $s_{\pi}=\Sigma_{\pi}^{+}v$ and kills the troublesome cross term in the second moment, and (ii) isolate the Edgeworth-gradient contribution in the lattice ratio lemma, yielding an $n^{-1}$-scale error term with polynomial growth in the standardized deviation $z=(N-\mathbb{E}[N])/\sqrt{n}$ whose moments are uniformly bounded.
Lemma 4.4 (Uniform moment bounds for standardized histograms).
Assume full support $\delta_{\mathrm{full}}>0$ and fix $\eta\in(0,1/2)$. Let $m\geq 1$ and $k$ satisfy $\eta m\leq k\leq(1-\eta)m$. Let $S$ be the histogram of $m$ shuffled messages under $T_{m,k}$, and define $z^{-}:=(S^{-}-\mathbb{E}[S]^{-})/\sqrt{m}\in\mathbb{R}^{d-1}$. Then for every integer $p\geq 1$ there exists $C_{p}=C_{p}(d,\delta_{\mathrm{full}},\eta)$ such that
$\mathbb{E}\bigl[\|z^{-}\|_{2}^{p}\bigr]\leq C_{p}.$
The same bound holds with $S$ replaced by $N$ (and $m$ replaced by $n$).
Proof.
Fix $y\in\mathcal{Y}$. Under $T_{m,k}$, the coordinate $S_{y}$ is a sum of $m$ independent Bernoulli variables with success probabilities in $\delta_{\mathrm{full}},1-\delta_{\mathrm{full}}]$ (exactly mkm-k draws from $W_{0}(y)$ and $k$ draws from $W_{1}(y)$). Bernstein’s inequality therefore gives constants $c_{0},c_{1}>0$ depending only on $\delta_{\mathrm{full}}$ such that for all $t\geq 0$,
$\mathbb{P}\bigl(|S_{y}-\mathbb{E}[S_{y}]|>t\bigr)\leq 2\exp\!\left(-\frac{c_{0}t^{2}}{m+c_{1}t}\right).$
Using the identity $\mathbb{E}[|X|^{p}]=\int_{0}^{\infty}pt^{p-1}\mathbb{P}(|X|>t)\,dt$ and splitting the integral at $t=m$ yields $\mathbb{E}[|S_{y}-\mathbb{E}[S_{y}]|^{p}]\leq C^{\prime}_{p}\,m^{p/2}$ for a constant $C^{\prime}_{p}=C^{\prime}_{p}(p,\delta_{\mathrm{full}})$. Since $d$ is fixed and $\|u\|_{2}\leq\sqrt{d}\,\|u\|_{\infty}$, a union bound over $y\leq d-1$ and Minkowski’s inequality imply $\mathbb{E}[\|S^{-}-\mathbb{E}[S]^{-}\|_{2}^{p}]\leq C^{\prime\prime}_{p}(d,\delta_{\mathrm{full}})\,m^{p/2}$. Dividing by $m^{p/2}$ gives the claim. ∎
Theorem 4.5 (Uniform sharpness for $k=\Theta(n)$).
Assume full support $\delta_{\mathrm{full}}>0$ and fix $\eta\in(0,1/2)$. Then, uniformly over all integers $k$ with $\eta n\leq k\leq(1-\eta)n$,
$\mathrm{JSD}(T_{n,k}\|T_{n,k+1})=\frac{1}{8n}I_{k/n}+O(n^{-2}),$
where $I_{\pi}=v^{\top}\Sigma_{\pi}^{+}v$ and $\Sigma_{\pi}=(1-\pi)\Sigma_{0}+\pi\Sigma_{1}$. The implicit constant depends only on $(d,\delta_{\mathrm{full}},\eta)$.
Proof.
Fix $n$ and $k\in\{\lceil\eta n\rceil,\dots,\lfloor(1-\eta)n\rfloor\}$. Let $P:=T_{n,k}$ and $Q:=T_{n,k+1}$. Write the likelihood ratio and its centered version as
$L:=\frac{dQ}{dP}=1+U,\qquad U=\mathbb{E}[r(Y_{\star})\mid N],$
where $r(y)=W_{1}(y)/W_{0}(y)-1$ and $Y_{\star}\sim W_{0}$ under $P$ (Lemma 3.3). We prove
$\mathbb{E}_{P}[D(1+U)]=\frac{1}{8n}I_{k/n}+O(n^{-2}),$ (9)
where $D$ is the pointwise JSD functional (Lemma 5.1).
Step 1: an exact regression identity and the leading $1/n$ second moment.
Let $\pi:=k/n\in[\eta,1-\eta]$ and $\Delta:=N-\mathbb{E}_{P}[N]$. First note $\mathbb{E}_{P}[U]=\mathbb{E}_{P}[L-1]=0$. Next, since $U=\mathbb{E}[r(Y_{\star})\mid N]$ and $\Delta$ is $N$-measurable,
$\mathbb{E}_{P}[\Delta\,U]=\mathbb{E}_{P}\bigl[\Delta\,\mathbb{E}[r(Y_{\star})\mid N]\bigr]=\mathbb{E}_{P}[\Delta\,r(Y_{\star})].$
Under $P$, we may write $N=e_{Y_{\star}}+S$ where $S$ is the histogram of the remaining n1n-1 users (with $k$ ones and nk1n-k-1 zeros), and $S$ is independent of $Y_{\star}$. Since $\mathbb{E}_{P}[r(Y_{\star})]=0$ and $\mathbb{E}_{P}[S-\mathbb{E}_{P}[S]]=0$, we have $\mathbb{E}_{P}[r(Y_{\star})(S-\mathbb{E}_{P}[S])]=0$. Therefore
$\mathbb{E}_{P}[\Delta\,r(Y_{\star})]=\mathbb{E}_{P}\bigl[(e_{Y_{\star}}-W_{0})\,r(Y_{\star})\bigr]=\sum_{y\in\mathcal{Y}}W_{0}(y)\,r(y)\,(e_{y}-W_{0}).$
Using $W_{0}(y)r(y)=W_{1}(y)-W_{0}(y)=v(y)$ and $\sum_{y}v(y)=0$, this equals $\sum_{y}v(y)e_{y}=v$. Hence
$\mathbb{E}_{P}[\Delta\,U]=v.$ (10)
Moreover, by fixed composition,
$\mathrm{Cov}_{P}(N)=\mathbb{E}_{P}[\Delta\Delta^{\top}]=(n-k)\Sigma_{0}+k\Sigma_{1}=n\Sigma_{\pi}.$ (11)
Let $s_{\pi}:=\Sigma_{\pi}^{+}v\in T$ and define the scalar score statistic
$Z:=s_{\pi}^{\top}\Delta.$
Define the residual
$R:=U-\frac{1}{n}Z.$
We claim $\mathbb{E}_{P}[RZ]=0$. Indeed, (10) and (11) imply
$\mathbb{E}_{P}[\Delta R]=\mathbb{E}_{P}[\Delta U]-\frac{1}{n}\mathbb{E}_{P}[\Delta\Delta^{\top}]s_{\pi}=v-\Sigma_{\pi}s_{\pi}=v-\Sigma_{\pi}\Sigma_{\pi}^{+}v=0,$
since $v\in T$ and $\Sigma_{\pi}\Sigma_{\pi}^{+}$ is the identity on $T$. Taking the inner product with $s_{\pi}$ yields $\mathbb{E}_{P}[RZ]=s_{\pi}^{\top}\mathbb{E}_{P}[\Delta R]=0$. Consequently,
$\mathbb{E}_{P}[U^{2}]=\mathbb{E}_{P}\!\left[\left(\frac{Z}{n}\right)^{\!2}\right]+\mathbb{E}_{P}[R^{2}].$ (12)
Finally, (11) gives
$\mathbb{E}_{P}[Z^{2}]=s_{\pi}^{\top}\mathrm{Cov}_{P}(N)\,s_{\pi}=n\,s_{\pi}^{\top}\Sigma_{\pi}s_{\pi}=n\,v^{\top}\Sigma_{\pi}^{+}v=nI_{\pi},$
hence
$\mathbb{E}_{P}\!\left[\left(\frac{Z}{n}\right)^{\!2}\right]=\frac{I_{\pi}}{n}.$ (13)
Thus, to prove (9) it suffices to show
$\sup_{\eta n\leq k\leq(1-\eta)n}\mathbb{E}_{P}[R^{2}]=O(n^{-2}),$ (14)
and to bound higher moments of $U$ at the corresponding scales.
Step 2: closing the “polynomial tail vs. exponential tail” gap in $L^{2}$.
We control $R$ in $L^{p}$ (hence in $L^{2}$) on an exponentially likely window using a refined ratio expansion that isolates the Edgeworth-gradient contribution (Lemma A.3 in Appendix A).
Let m:=n1m:=n-1 and $\pi_{m}:=k/m$. Let $S$ be the histogram of the remaining $m$ users, so $S$ has law $T_{m,k}$ under $P$. For $y\in\mathcal{Y}$ write $p_{y}(N):=\mathbb{P}(S=N-e_{y})$ (with the convention $p_{y}(N)=0$ if $N_{y}=0$). Bayes’ rule (cf. (7)) gives
$\mathbb{P}(Y_{\star}=y\mid N)=\frac{W_{0}(y)\,p_{y}(N)}{\sum_{y^{\prime}\in\mathcal{Y}}W_{0}(y^{\prime})\,p_{y^{\prime}}(N)}.$ (15)
Fix the baseline symbol $d$ and write $s:=N-e_{d}\in\mathbb{Z}^{d}$ (so $\mathbf{1}^{\top}s=m$). For $y\in\mathcal{Y}$ define the shift $t_{y}:=e_{d}-e_{y}\in\mathbb{Z}^{d}$, so $s-t_{y}=N-e_{y}$ and $\mathbf{1}^{\top}t_{y}=0$, $\|t_{y}\|_{1}\leq 2$.
Let $\mu:=\mathbb{E}[S]=(m-k)W_{0}+kW_{1}$ and $\Sigma_{\mathrm{ps}}:=\Sigma_{\pi_{m}}=(1-\pi_{m})\Sigma_{0}+\pi_{m}\Sigma_{1}$ be the per-sample covariance for the $m$-sum. Define the standardized deviation (in reduced coordinates as in Appendix A)
$z^{-}:=\frac{s^{-}-\mu^{-}}{\sqrt{m}}\in\mathbb{R}^{d-1}.$
Lemma A.3 gives, for each $y\in\mathcal{Y}$,
$\log\frac{p_{y}(N)}{p_{d}(N)}=\frac{1}{m}\,t_{y}^{\top}\Sigma_{\mathrm{ps}}^{+}(s-\mu)-\frac{1}{2m}\,t_{y}^{\top}\Sigma_{\mathrm{ps}}^{+}t_{y}+\frac{1}{m}\,\mathcal{Q}_{m,t_{y}}(z^{-})+\rho_{m,t_{y}}(s),$ (16)
where $\mathcal{Q}_{m,t_{y}}$ is a quadratic polynomial with coefficients bounded by a constant depending only on $(d,\delta_{\mathrm{full}},\eta)$, and the remainder satisfies
$|\rho_{m,t_{y}}(s)|\leq C\,m^{-3/2}\bigl(1+\|z^{-}\|_{2}^{9}\bigr).$ (17)
Introduce the $T$-valued tilt vector
$\theta:=\frac{1}{m}\,\Sigma_{\mathrm{ps}}^{+}(s-\mu)\in T.$
Rewrite (16) as
$\log\frac{p_{y}(N)}{p_{d}(N)}=t_{y}^{\top}\theta+\gamma_{y}+\frac{1}{m}\,\mathcal{Q}_{m,t_{y}}(z^{-})+\rho_{m,t_{y}}(s),$ (18)
where $\gamma_{y}:=-(2m)^{-1}t_{y}^{\top}\Sigma_{\mathrm{ps}}^{+}t_{y}$ is deterministic and uniformly $O(m^{-1})$. Substituting (18) into (15) cancels the common factor $p_{d}(N)$ and yields
$\mathbb{P}(Y_{\star}=y\mid N)=\frac{W_{0}(y)\exp(t_{y}^{\top}\theta)\,\exp(\gamma_{y}+(1/m)\mathcal{Q}_{m,t_{y}}(z^{-})+\rho_{m,t_{y}}(s))}{\sum_{y^{\prime}}W_{0}(y^{\prime})\exp(t_{y^{\prime}}^{\top}\theta)\,\exp(\gamma_{y^{\prime}}+(1/m)\mathcal{Q}_{m,t_{y^{\prime}}}(z^{-})+\rho_{m,t_{y^{\prime}}}(s))}.$
Because $t_{y}=e_{d}-e_{y}$, the map $y\mapsto t_{y}^{\top}\theta$ is of the form $c(\theta)-e_{y}^{\top}\theta$; the common term cancels between numerator and denominator. Since $\|\theta\|=O_{P}(m^{-1/2})$ and $\gamma_{y}+(1/m)\mathcal{Q}_{m,t_{y}}(z^{-})+\rho_{m,t_{y}}(s)=O_{P}(m^{-1})$, a Taylor expansion of both numerator and normalizer gives, uniformly in $y$,
$\mathbb{P}(Y_{\star}=y\mid N)=W_{0}(y)\Bigl(1+(e_{y}-W_{0})^{\top}\theta\Bigr)+\varepsilon_{y}(N),$ (19)
where the error terms satisfy
$|\varepsilon_{y}(N)|\leq C\left(\frac{1+\|z^{-}\|_{2}^{2}}{m}+\frac{1+\|z^{-}\|_{2}^{9}}{m^{3/2}}\right)$ (20)
for a constant $C=C(d,\delta_{\mathrm{full}},\eta)$.
Compute $U=\mathbb{E}[r(Y_{\star})\mid N]=\sum_{y}r(y)\mathbb{P}(Y_{\star}=y\mid N)$ using (19). The constant term $\sum_{y}r(y)W_{0}(y)=\sum_{y}v(y)=0$ vanishes. For the linear term, $r(y)W_{0}(y)=v(y)$ and $\sum_{y}v(y)W_{0}=0$, hence
$\sum_{y}r(y)W_{0}(y)\,(e_{y}-W_{0})=\sum_{y}v(y)e_{y}=v.$
Therefore (19) yields
$U=v^{\top}\theta+\varepsilon(N),\qquad|\varepsilon(N)|\leq C\left(\frac{1+\|z^{-}\|_{2}^{2}}{m}+\frac{1+\|z^{-}\|_{2}^{9}}{m^{3/2}}\right),$ (21)
where $\varepsilon(N):=\sum_{y}r(y)\varepsilon_{y}(N)$ and we used boundedness of $r$ under full support. Since $\theta=\frac{1}{m}\Sigma_{\mathrm{ps}}^{+}(s-\mu)$ and $s-\mu=(N-\mathbb{E}_{P}[N])+(e_{d}-W_{0})$, we have
$v^{\top}\theta=\frac{1}{m}\,v^{\top}\Sigma_{\mathrm{ps}}^{+}(N-\mathbb{E}_{P}[N])+O(m^{-1}).$
Because $\pi_{m}-\pi=O(1/n)$ and $\Sigma_{\pi}$ is affine in $\pi$, the pseudoinverse map $\pi\mapsto\Sigma_{\pi}^{+}$ is Lipschitz on $\eta/2,1-\eta/2]$; therefore $v^{\top}\Sigma_{\mathrm{ps}}^{+}(N-\mathbb{E}_{P}[N])=v^{\top}\Sigma_{\pi}^{+}(N-\mathbb{E}_{P}[N])+O_{P}(n^{-1/2})$. Combining with $m^{-1}=n^{-1}+O(n^{-2})$ yields
$v^{\top}\theta=\frac{1}{n}\,s_{\pi}^{\top}\Delta+O_{P}(n^{-3/2})+O(n^{-1}).$
Absorb the deterministic $O(n^{-1})$ term into (21) to obtain
$R=U-\frac{1}{n}Z=O_{P}\!\left(\frac{1+\|z^{-}\|_{2}^{2}}{n}\right)+O_{P}\!\left(\frac{1+\|z^{-}\|_{2}^{9}}{n^{3/2}}\right).$ (22)
Finally, uniform moment bounds for $z^{-}$ follow from bounded increments and fixed $d$: for every fixed $p\geq 1$,
$\sup_{\eta n\leq k\leq(1-\eta)n}\mathbb{E}_{P}[\|z^{-}\|_{2}^{p}]\leq C_{p}(d,\delta_{\mathrm{full}},\eta)<\infty.$
Applying this with $p=20$ and squaring (22) gives (14) and moreover $\mathbb{E}_{P}[R^{4}]=O(n^{-4})$ uniformly in $k$.
Step 3: Taylor expansion for JSD and completion.
Let $B:=\sup_{y}|r(y)|<\infty$ (bounded under full support), so $|U|\leq B$ almost surely. Define the event $G_{n}:=\{|U|\leq 1/2\}$. Since $U=Z/n+R$, Bernstein’s inequality for the sum $Z$ (a sum of $n$ bounded independent increments) gives $\mathbb{P}(|Z|>n/4)\leq 2e^{-cn}$, and Markov’s inequality with $\mathbb{E}_{P}[R^{4}]=O(n^{-4})$ gives $\mathbb{P}(|R|>1/4)=O(n^{-4})$. Therefore $\mathbb{P}(G_{n}^{c})\leq 2e^{-cn}+O(n^{-4})$, uniformly in $k$.
On $G_{n}$ apply Lemma 5.2 (with $u=U$):
$D(1+U)=\frac{U^{2}}{8}-\frac{U^{3}}{16}+\frac{7U^{4}}{192}-\frac{3U^{5}}{128}+R_{6}(U),\qquad|R_{6}(U)|\leq C_{6}|U|^{6}.$
Taking expectations and using boundedness of $D$ on $[0,1+B]$ yields
$\mathbb{E}_{P}[D(1+U)]=\frac{1}{8}\mathbb{E}_{P}[U^{2}]-\frac{1}{16}\mathbb{E}_{P}[U^{3}]+\frac{7}{192}\mathbb{E}_{P}[U^{4}]+O\!\left(\mathbb{E}_{P}[|U|^{5}]\right)+O(e^{-cn}).$ (23)
Using $U=Z/n+R$, Rosenthal moment bounds for $Z$ (so $\mathbb{E}[Z^{4}]=O(n^{2})$ and $\mathbb{E}[|Z|^{5}]=O(n^{5/2})$), and $\mathbb{E}[R^{p}]=O(n^{-p})$ for $p\in\{2,4\}$, we obtain
$\mathbb{E}_{P}[U^{4}]=O(n^{-2}),\qquad\mathbb{E}_{P}[|U|^{5}]=O(n^{-5/2}),$
uniformly in $k$. For the cubic moment, expand and use $\mathbb{E}[Z^{3}]=O(n)$:
$\mathbb{E}_{P}[U^{3}]=\frac{1}{n^{3}}\mathbb{E}_{P}[Z^{3}]+\frac{3}{n^{2}}\mathbb{E}_{P}[Z^{2}R]+\frac{3}{n}\mathbb{E}_{P}[ZR^{2}]+\mathbb{E}_{P}[R^{3}].$
By Cauchy–Schwarz, $|\mathbb{E}[Z^{2}R]|\leq(\mathbb{E}[Z^{4}])^{1/2}(\mathbb{E}[R^{2}])^{1/2}=O(n)\cdot O(n^{-1})=O(1)$ and $|\mathbb{E}[ZR^{2}]|\leq(\mathbb{E}[Z^{2}])^{1/2}(\mathbb{E}[R^{4}])^{1/2}=O(\sqrt{n})\cdot O(n^{-2})=O(n^{-3/2})$, while $|\mathbb{E}[R^{3}]|\leq(\mathbb{E}[R^{4}])^{3/4}=O(n^{-3})$. Therefore $\mathbb{E}_{P}[U^{3}]=O(n^{-2})$ uniformly in $k$.
Substitute these bounds into (23) and use (12)–(13) plus (14):
$\mathbb{E}_{P}[D(1+U)]=\frac{1}{8}\left(\frac{I_{\pi}}{n}+\mathbb{E}_{P}[R^{2}]\right)+O(n^{-2})=\frac{I_{k/n}}{8n}+O(n^{-2}),$
uniformly over $\eta n\leq k\leq(1-\eta)n$. This proves (9), i.e. the stated uniform JSD expansion. ∎
5 Sharp Jensen–Shannon expansions
5.1 A pointwise functional representation
Lemma 5.1 (JSD as an expectation of a pointwise functional).
Let $P,Q$ be distributions on a finite space with $Q\ll P$ and likelihood ratio $L=dQ/dP$. Then
$\mathrm{JSD}(P\|Q)=\mathbb{E}_{P}[D(L)],$
where for $t\geq 0$,
$D(t):=\frac{1}{2}\log\frac{2}{1+t}+\frac{t}{2}\log\frac{2t}{1+t},\qquad D(0):=\frac{\log 2}{2}.$
Proof.
Let $M=(P+Q)/2$. Then
$\mathrm{JSD}(P\|Q)=\frac{1}{2}\sum_{x}P(x)\log\frac{P(x)}{M(x)}+\frac{1}{2}\sum_{x}Q(x)\log\frac{Q(x)}{M(x)}.$
Using $Q=LP$ and $M=\frac{1}{2}(1+L)P$ yields the stated integrand after rearrangement. ∎
5.2 Canonical neighboring pair: sharp $O(n^{-3})$ expansion
Assume $\delta_{\star}>0$ and define (with $Y\sim W_{0}$)
$\chi^{2}:=\chi^{2}(W_{1}\|W_{0})=\mathbb{E}[r(Y)^{2}],\qquad\mu_{3}:=\mathbb{E}[r(Y)^{3}].$
Lemma 5.2 (Sixth-order Taylor expansion).
For $|u|\leq\tfrac{1}{2}$,
$D(1+u)=\frac{u^{2}}{8}-\frac{u^{3}}{16}+\frac{7u^{4}}{192}-\frac{3u^{5}}{128}+R_{6}(u),\qquad|R_{6}(u)|\leq C_{6}|u|^{6},$
for a finite constant $C_{6}$.
Proof.
$D$ is analytic on $(0,\infty)$. Expand around $u=0$ and apply Taylor’s theorem with Lagrange remainder on the compact interval $[1/2,3/2]$. ∎
Lemma 5.3 (Moments of the score average).
Let $Z=r(Y)$ for $Y\sim W_{0}$ and $U_{n}=n^{-1}\sum_{i=1}^{n}Z_{i}$ with i.i.d. copies. Then
$\mathbb{E}[U_{n}^{2}]=\frac{\chi^{2}}{n},\qquad\mathbb{E}[U_{n}^{3}]=\frac{\mu_{3}}{n^{2}},$
and moreover
$\mathbb{E}[U_{n}^{4}]=\frac{3(\chi^{2})^{2}}{n^{2}}+O(n^{-3}),\qquad\mathbb{E}[|U_{n}|^{6}]=O(n^{-3}).$
Proof.
Write $S_{n}=\sum_{i=1}^{n}Z_{i}$, so that $U_{n}=S_{n}/n$. Since $\mathbb{E}[Z]=0$ and $\mathrm{Var}(Z)=\chi^{2}$, we have $\mathbb{E}[S_{n}^{2}]=n\chi^{2}$, hence $\mathbb{E}[U_{n}^{2}]=\chi^{2}/n$. For the third moment, expand $S_{n}^{3}=\sum_{i}Z_{i}^{3}+3\sum_{i\neq j}Z_{i}^{2}Z_{j}+6\sum_{i<j<k}Z_{i}Z_{j}Z_{k}$ and use independence and centering to get $\mathbb{E}[S_{n}^{3}]=n\mathbb{E}[Z^{3}]=n\mu_{3}$, hence $\mathbb{E}[U_{n}^{3}]=\mu_{3}/n^{2}$.
For the fourth moment, independence and centering give
$\mathbb{E}[S_{n}^{4}]=\sum_{i=1}^{n}\mathbb{E}[Z_{i}^{4}]+6\sum_{1\leq i<j\leq n}\mathbb{E}[Z_{i}^{2}]\mathbb{E}[Z_{j}^{2}]=n\mathbb{E}[Z^{4}]+3n(n-1)(\chi^{2})^{2}.$
Therefore $\mathbb{E}[U_{n}^{4}]=3(\chi^{2})^{2}/n^{2}+O(n^{-3})$, since $\mathbb{E}[Z^{4}]<\infty$ (recall $|Z|\leq\delta_{\star}^{-1}$).
Finally, for the sixth absolute moment, Rosenthal’s inequality (for $p=6$; see, e.g., [21, Chapter 2, Section 9]) yields
$\mathbb{E}[|S_{n}|^{6}]\leq C\Bigl(n\,\mathbb{E}[|Z|^{6}]+(n\,\mathrm{Var}(Z))^{3}\Bigr)=O(n^{3}),$
because $Z\in[-1,\delta_{\star}^{-1}-1]$ is bounded under $\delta_{\star}>0$. Dividing by $n^{6}$ gives $\mathbb{E}[|U_{n}|^{6}]=O(n^{-3})$. ∎
Theorem 5.4 (Sharp canonical JSD expansion).
Assume $\delta_{\star}>0$ and consider $(T_{n,0},T_{n,1})$. Then, as $n\to\infty$,
$\mathrm{JSD}(T_{n,0}\|T_{n,1})=\frac{\chi^{2}}{8n}-\frac{\mu_{3}}{16n^{2}}+\frac{7}{64}\frac{(\chi^{2})^{2}}{n^{2}}+O(n^{-3}),$
where the implicit constant depends only on $\delta_{\star}$ (through uniform moment bounds).
Proof.
By Lemmas 3.2 and 5.1, $\mathrm{JSD}(T_{n,0}\|T_{n,1})=\mathbb{E}[D(1+U_{n})]$. Let $B_{n}=\{|U_{n}|\leq 1/2\}$. Since $Z_{i}\in[-1,\delta_{\star}^{-1}-1]$, Hoeffding’s inequality [15] implies $\mathbb{P}(B_{n}^{c})\leq 2e^{-cn}$ for a constant $c(\delta_{\star})>0$. On $B_{n}$ apply Lemma 5.2 and take expectations:
$\mathbb{E}[D(1+U_{n})\mathbf{1}_{B_{n}}]=\frac{1}{8}\mathbb{E}[U_{n}^{2}\mathbf{1}_{B_{n}}]-\frac{1}{16}\mathbb{E}[U_{n}^{3}\mathbf{1}_{B_{n}}]+\frac{7}{192}\mathbb{E}[U_{n}^{4}\mathbf{1}_{B_{n}}]-\frac{3}{128}\mathbb{E}[U_{n}^{5}\mathbf{1}_{B_{n}}]+\mathbb{E}[R_{6}(U_{n})\mathbf{1}_{B_{n}}].$
Since $U_{n}$ is bounded ($|U_{n}|\leq\delta_{\star}^{-1}$), each moment satisfies $|\mathbb{E}[U_{n}^{k}\mathbf{1}_{B_{n}}]-\mathbb{E}[U_{n}^{k}]|\leq\mathbb{E}[|U_{n}|^{k}\mathbf{1}_{B_{n}^{c}}]\leq\delta_{\star}^{-k}\cdot 2e^{-cn}$, so we may replace $\mathbb{E}[U_{n}^{k}\mathbf{1}_{B_{n}}]$ by $\mathbb{E}[U_{n}^{k}]$ at exponentially negligible cost. Using Lemma 5.3, the first three terms yield $\chi^{2}/(8n)-\mu_{3}/(16n^{2})+(7/64)(\chi^{2})^{2}/n^{2}+O(n^{-3})$. The $U_{n}^{5}$ and remainder terms are $O(n^{-3})$ by moment bounds; the tail $\mathbb{E}[D(1+U_{n})\mathbf{1}_{B_{n}^{c}}]$ is exponentially small since $D$ is bounded on $0,\delta_{\star}^{-1}]$. ∎
5.3 Proportional compositions: universal leading constant
Assume full support $\delta_{\mathrm{full}}>0$ and let $k=k(n)$ with $k/n\to\pi\in(0,1)$.
Theorem 5.5 (Leading JSD constant for $k/n\to\pi$).
Assume $\delta_{\mathrm{full}}>0$ and $k/n\to\pi\in(0,1)$. Then
$\mathrm{JSD}(T_{n,k}\|T_{n,k+1})=\frac{I_{\pi}}{8n}+o(n^{-1}),\qquad I_{\pi}=v^{\top}\Sigma_{\pi}^{+}v,$
with $\Sigma_{\pi}$ as in (1).
Proof.
Let $P=T_{n,k}$ and $Q=T_{n,k+1}$ and write $L=dQ/dP=1+U$ with $U=\mathbb{E}[r(Y_{\star})\mid N]$ by Lemma 3.3. By Theorem 4.1, on $A_{n}$,
$U=\frac{1}{n}s_{n}^{\top}(N-\mathbb{E}[N])+R_{n}(N),\qquad\sup_{A_{n}}|R_{n}|\leq Cn^{-3/4},$
where $s_{n}=\Sigma_{\pi_{n}}^{+}v$ and $\pi_{n}=k/(n-1)\to\pi$. In particular $U=O_{P}(n^{-1/2})$ and $|U|\leq C$ almost surely by full support. Taylor expanding $D(1+u)=u^{2}/8+O(|u|^{3})$ and using Lemma 5.1 yields
$\mathrm{JSD}(P\|Q)=\frac{1}{8}\mathbb{E}[U^{2}]+\mathbb{E}[O(|U|^{3})].$
To bound the cubic term: the leading part of $U$ is $(1/n)s_{n}^{\top}(N-\mathbb{E}[N])=O_{P}(n^{-1/2})$, hence $\mathbb{E}[|U|^{3}]=O(n^{-3/2})=o(n^{-1})$ (using boundedness of $U$ and Rosenthal’s inequality for the leading sum). The contribution from $A_{n}^{c}$ is exponentially negligible. Therefore $\mathrm{JSD}(P\|Q)=\frac{1}{8}\mathbb{E}[U^{2}]+o(n^{-1})$. Next,
$\mathbb{E}[U^{2}]=\frac{1}{n^{2}}\mathbb{E}\bigl[(s_{n}^{\top}(N-\mathbb{E}[N]))^{2}\bigr]+o(n^{-1}).$
Under fixed composition with $k$ ones and nkn-k zeros, $\mathrm{Cov}(N)=(n-k)\Sigma_{0}+k\Sigma_{1}=n\Sigma_{k/n}$ exactly. The $o(n)$ remainder arises only from replacing $k/n$ by $\pi$ and $s_{n}$ by $s=\Sigma_{\pi}^{+}v$. Hence
$\frac{1}{n}\mathbb{E}\bigl[(s_{n}^{\top}(N-\mathbb{E}[N]))^{2}\bigr]\to s^{\top}\Sigma_{\pi}s=v^{\top}\Sigma_{\pi}^{+}v=I_{\pi},$
and so $\mathbb{E}[U^{2}]=I_{\pi}/n+o(n^{-1})$, proving the claim. ∎
Theorem 5.6 (Universal smooth $f$-divergence constant).
Let $f$ be three-times continuously differentiable near $1$ with $f(1)=f^{\prime}(1)=0$. Under the assumptions of Theorem 5.5,
$D_{f}(T_{n,k+1}\|T_{n,k})=\mathbb{E}_{T_{n,k}}[f(L_{n,k})]=\frac{f^{\prime\prime}(1)}{2}\frac{I_{\pi}}{n}+o(n^{-1}).$
In particular, for fixed $\alpha>1$,
$D_{\alpha}(T_{n,k+1}\|T_{n,k})=\frac{\alpha}{2n}I_{\pi}+o(n^{-1}).$
Proof.
For a general $f$-divergence, Taylor expand $f(1+u)=\frac{1}{2}f^{\prime\prime}(1)u^{2}+O(|u|^{3})$ and use $\mathbb{E}[U^{2}]=I_{\pi}/n+o(n^{-1})$ and $\mathbb{E}|U|^{3}=o(n^{-1})$ (boundedness plus $U=O_{P}(n^{-1/2})$). For Rényi divergence, note that the $f$-divergence with $f(t)=(t^{\alpha}-1)/(\alpha-1)$ (which has $f^{\prime\prime}(1)=\alpha$) gives $D_{f}=\frac{\alpha}{2}\frac{I_{\pi}}{n}+o(n^{-1})$, and $D_{\alpha}=\frac{1}{\alpha-1}\log\bigl(1+(\alpha-1)D_{f}\bigr)=D_{f}+O(D_{f}^{2})=\frac{\alpha}{2n}I_{\pi}+o(n^{-1})$, since $D_{f}=O(n^{-1})$. ∎
6 Gaussian Differential Privacy equivalence
We now show that the privacy loss is asymptotically Gaussian under both hypotheses with explicit $O(n^{-1/2})$ bounds, yielding equivalence to GDP and the full limiting $(\varepsilon,\delta)$ curve.
6.1 Berry–Esseen under $T_{n,0}$ (canonical) and under $T_{n,k}$ (general)
Lemma 6.1 (Berry–Esseen for i.i.d. averages).
Let $Z_{1},Z_{2},\dots$ be i.i.d. with $\mathbb{E}[Z_{1}]=0$, $\mathrm{Var}(Z_{1})=\sigma^{2}>0$, and $\rho_{3}=\mathbb{E}|Z_{1}|^{3}<\infty$. Let $U_{n}=n^{-1}\sum_{i=1}^{n}Z_{i}$ and $\sigma_{n}=\sigma/\sqrt{n}$. Then (see, e.g., Petrov [21] or Serfling [22])
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{n}}{\sigma_{n}}\leq t\right)-\Phi(t)\right|\leq C_{\mathrm{BE}}\frac{\rho_{3}}{\sigma^{3}\sqrt{n}}$
for an absolute constant $C_{\mathrm{BE}}$.
6.2 Normal approximation for $\Lambda_{n}=\log(1+U_{n})$ under $T_{n,0}$
Assume $\delta_{\mathrm{full}}>0$ so that $U_{n}>-1$ and $\Lambda_{n}$ is bounded. Let $\sigma^{2}=\chi^{2}(W_{1}\|W_{0})$ and $\sigma_{n}=\sigma/\sqrt{n}$. Define $m_{n}^{-}=-\sigma_{n}^{2}/2=-\chi^{2}/(2n)$.
Theorem 6.2 (Normal approximation under $T_{n,0}$).
Assume $\delta_{\mathrm{full}}>0$. There exists $C<\infty$ (depending only on $\delta_{\mathrm{full}}$ and $\rho_{3}/\sigma^{3}$) such that, for all $n$,
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}_{T_{n,0}}\!\left(\frac{\Lambda_{n}-m_{n}^{-}}{\sigma_{n}}\leq t\right)-\Phi(t)\right|\leq\frac{C}{\sqrt{n}}.$
Moreover $\mathbb{E}_{T_{n,0}}[\Lambda_{n}]=m_{n}^{-}+O(n^{-3/2})$ and $\mathrm{Var}_{T_{n,0}}(\Lambda_{n})=\sigma_{n}^{2}+O(n^{-3/2})$.
Proof.
Write $\{\Lambda_{n}\leq m_{n}^{-}+\sigma_{n}t\}=\{U_{n}\leq a_{n}(t)\}$ where $a_{n}(t)=e^{m_{n}^{-}+\sigma_{n}t}-1$. A Taylor expansion gives $a_{n}(t)/\sigma_{n}=t+O(\sigma_{n}(1+t^{2}))$. Combine Lemma 6.1 with a Lipschitz bound for $\Phi$ and split into $|t|\leq n^{1/8}$ and its complement (Gaussian tails) to obtain the uniform $O(n^{-1/2})$ bound. Mean and variance follow from $\log(1+u)=u-\frac{1}{2}u^{2}+O(|u|^{3})$ and moment rates for $U_{n}$. ∎
6.3 Normal approximation under the alternative via change of measure
Let $T_{n,1}$ denote the alternative for the canonical pair and note the identity
$\mathbb{E}_{T_{n,1}}[f]=\mathbb{E}_{T_{n,0}}[(1+U_{n})f],$
from Lemma 3.2. Define $m_{n}^{+}=+\sigma_{n}^{2}/2=+\chi^{2}/(2n)$.
Theorem 6.3 (Normal approximation under $T_{n,1}$).
Assume $\delta_{\mathrm{full}}>0$. There exists $C<\infty$ such that, for all $n$,
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}_{T_{n,1}}\!\left(\frac{\Lambda_{n}-m_{n}^{+}}{\sigma_{n}}\leq t\right)-\Phi(t)\right|\leq\frac{C}{\sqrt{n}}.$
Moreover $\mathbb{E}_{T_{n,1}}[\Lambda_{n}]=m_{n}^{+}+O(n^{-3/2})$ and $\mathrm{Var}_{T_{n,1}}(\Lambda_{n})=\sigma_{n}^{2}+O(n^{-3/2})$.
Proof.
Fix $t$ and write $\{\Lambda_{n}\leq m_{n}^{+}+\sigma_{n}t\}=\{U_{n}\leq a_{n}(t)\}$ with $a_{n}(t)=e^{m_{n}^{+}+\sigma_{n}t}-1$. By change of measure,
$\mathbb{P}_{T_{n,1}}(\Lambda_{n}\leq m_{n}^{+}+\sigma_{n}t)=\mathbb{E}_{T_{n,0}}\bigl[(1+U_{n})\mathbf{1}\{U_{n}\leq a_{n}(t)\}\bigr]=\mathbb{P}(U_{n}\leq a_{n}(t))+\mathbb{E}[U_{n}\mathbf{1}\{U_{n}\leq a_{n}(t)\}].$
The first term is treated as in Theorem 6.2. For the second term, use that $U_{n}$ is almost surely bounded: $U_{n}\in[-1,B]$ with $B=\delta_{\star}^{-1}-1$. Define the compactly supported truncation
$g_{a,B}(u):=\begin{cases}0,&u\leq-1,\\ u,&-1<u\leq\min\{a,B\},\\ 0,&u>\min\{a,B\}.\end{cases}$
Then $g_{a,B}(U_{n})=U_{n}\mathbf{1}\{U_{n}\leq a\}$ almost surely and $g_{a,B}$ has bounded variation uniformly in $a$ (total variation $\leq 2B+2$). By integration by parts (Stieltjes form: $\mathbb{E}[g(X)]=g(b^{-})\mathbb{P}(X\leq b)-\int g^{\prime}(u)\mathbb{P}(X\leq u)\,du$ for $g$ of bounded variation and $X$ supported on [1,B][-1,B]; see, e.g., [12, Theorem 1, Chapter XVI]), the expectation of $g_{a,B}(U_{n})$ differs from that of $g_{a,B}(\sigma_{n}G)$ (with $G\sim N(0,1)$) by at most $\mathrm{TV}(g_{a,B})\cdot\sup_{u}|\mathbb{P}(U_{n}\leq u)-\Phi(u/\sigma_{n})|=O(n^{-1/2})$ uniformly in $a$, with the bound controlled by the Kolmogorov distance in Lemma 6.1. Since $g_{a,B}$ agrees with $u\mathbf{1}\{u\leq a\}$ on [1,B][-1,B] and $\sigma_{n}G$ exceeds $B$ only with exponentially small probability, we obtain a uniform Gaussian approximation for $\mathbb{E}[U_{n}\mathbf{1}\{U_{n}\leq a_{n}(t)\}]$ with $O(n^{-1/2})$ error. Combining the two terms yields the stated $O(n^{-1/2})$ bound; mean and variance follow by expanding $(1+u)\log(1+u)$ and $(1+u)\log^{2}(1+u)$ and using moment bounds. ∎
6.4 GDP and the full privacy curve
The Gaussian shift experiment $G_{\mu}=(N(0,1),N(\mu,1))$ defines $\mu$-GDP [7]. Conversions to $(\varepsilon,\delta)$ use the explicit GDP curve
$\delta_{\mathrm{GDP}}(\varepsilon;\mu)=\Phi\!\left(-\frac{\varepsilon}{\mu}+\frac{\mu}{2}\right)-e^{\varepsilon}\Phi\!\left(-\frac{\varepsilon}{\mu}-\frac{\mu}{2}\right).$ (24)
Theorem 6.4 (Asymptotic GDP equivalence and trade-off transfer).
Assume $\delta_{\mathrm{full}}>0$. For the canonical pair $(T_{n,0},T_{n,1})$, let $\mu_{n}=\sqrt{\chi^{2}(W_{1}\|W_{0})/n}$. Then the trade-off function $\mathcal{T}$ satisfies
$\mathcal{T}_{T_{n,0},T_{n,1}}(\alpha)=\mathcal{T}_{\mu_{n}}(\alpha)+O(n^{-1/2})$
uniformly in $\alpha\in[0,1]$, where $\mathcal{T}_{\mu}$ is the Gaussian shift trade-off. The same statement holds for $(T_{n,k},T_{n,k+1})$ with $\mu_{n}(\pi)=\sqrt{I_{\pi}/n}$ when $k/n\to\pi\in(0,1)$.
Proof.
By Neyman–Pearson, the optimal test is a threshold on the privacy loss. For the canonical pair, Theorems 6.2 and 6.3 give uniform $O(n^{-1/2})$ control of the privacy loss CDFs under both hypotheses with means $\mp\mu_{n}^{2}/2$ and variance $\mu_{n}^{2}$. This yields uniform closeness of the induced trade-off curve to the Gaussian trade-off.
For proportional compositions ($k/n\to\pi$), the argument proceeds in three steps: (i) Theorem 4.1 linearizes the log-likelihood on a typical set; (ii) a triangular-array Berry–Esseen bound (see, e.g., [21, 22]) for the score statistic $s^{\top}(N-\mathbb{E}[N])$ gives $O(n^{-1/2})$ normal approximation under $T_{n,k}$; (iii) the alternative-hypothesis approximation under $T_{n,k+1}$ follows by contiguity (Remark 6.5 below). The full LAN expansion making these steps precise is Theorem 7.3. ∎
Remark 6.5 (Normal approximation under both hypotheses for general $k$).
For the canonical pair ($k=0$), the normal approximation under the alternative $T_{n,1}$ is proved explicitly in Theorem 6.3 via change of measure. For general $(T_{n,k},T_{n,k+1})$ with $k/n\to\pi$, the corresponding result under $T_{n,k+1}$ follows by contiguity: since $\Lambda_{n,k}=\log L_{n,k}$ is asymptotically $N(-\mu^{2}/2,\mu^{2})$ under $T_{n,k}$ (with $\mu=\mu_{n}(\pi)\to 0$), Le Cam’s third lemma implies that under $T_{n,k+1}$, $\Lambda_{n,k}\Rightarrow N(+\mu^{2}/2,\mu^{2})$. The $O(n^{-1/2})$ rate transfers via the Berry–Esseen bound and the bounded-variation integration-by-parts argument used in the proof of Theorem 6.3 (specifically, the Stieltjes transfer from the Kolmogorov distance to expectations of BV functions, applied to the change-of-measure weight $(1+U_{n,k})\mathbf{1}\{U_{n,k}\leq a\}$).
Corollary 6.6 (Asymptotic $(\varepsilon,\delta)$ curve).
Under the assumptions of Theorem 6.4, for each fixed $\varepsilon\geq 0$,
$\delta_{T_{n,k+1}\|T_{n,k}}(\varepsilon)=\delta_{\mathrm{GDP}}(\varepsilon;\mu_{n}(\pi))+O(n^{-1/2})$
uniformly for $\varepsilon$ in compact subsets of $0,\infty)$.
7 Local Asymptotic Normality and Le Cam equivalence
We formulate Local Asymptotic Normality (LAN) for the shuffle model; for background on LAN theory see van der Vaart [27] and Le Cam–Yang [19].
7.1 Trade-off and Le Cam distance for binary experiments
For a binary experiment $E=(P,Q)$ define its trade-off function
$\mathcal{T}_{P,Q}(\alpha):=\inf\{\mathbb{E}_{Q}[1-\phi]:\phi:\Omega\to[0,1],\ \mathbb{E}_{P}[\phi]\leq\alpha\},\qquad\alpha\in[0,1].$
Definition 7.1 (Deficiency and Le Cam distance).
Let $E=(P,Q)$ and $F=(P^{\prime},Q^{\prime})$ be binary experiments (possibly on different spaces). The deficiency is
$\delta(E,F):=\inf_{K}\max\{\|KP-P^{\prime}\|_{\mathrm{TV}},\|KQ-Q^{\prime}\|_{\mathrm{TV}}\},$
where the infimum ranges over Markov kernels $K$ mapping the sample space of $E$ to that of $F$. The Le Cam distance is $\Delta(E,F):=\max\{\delta(E,F),\delta(F,E)\}$.
Lemma 7.2 (Trade-off control implies Le Cam control).
If $\sup_{\alpha\in[0,1]}|\mathcal{T}_{E}(\alpha)-\mathcal{T}_{F}(\alpha)|\leq\eta$, then $\Delta(E,F)\leq\eta$.
Proof.
For a binary experiment $E=(P,Q)$ define the testing risk set
$\mathcal{R}(E):=\left\{(\alpha,\beta)\in[0,1]^{2}:\ \exists\ \text{test }\phi\in[0,1]\ \text{ with }\ \alpha=\mathbb{E}_{P}[\phi],\ \beta=\mathbb{E}_{Q}[1-\phi]\right\}.$
Randomization of tests implies that $\mathcal{R}(E)$ is convex and closed, and by definition $\mathcal{T}_{E}(\alpha)=\inf\{\beta:(\alpha,\beta)\in\mathcal{R}(E)\}$. Thus the uniform bound $\sup_{\alpha}|\mathcal{T}_{E}(\alpha)-\mathcal{T}_{F}(\alpha)|\leq\eta$ implies mutual $\eta$-containment of testing risk sets:
$\mathcal{R}(E)\subset\mathcal{R}(F)+[0,\eta]^{2},\qquad\mathcal{R}(F)\subset\mathcal{R}(E)+[0,\eta]^{2}.$
By the randomization theorem for binary experiments (equivalently, the Blackwell–Le Cam ordering; see, e.g., [19, Ch. 2] or [26]), these inclusions are equivalent to $\delta(E,F)\leq\eta$ and $\delta(F,E)\leq\eta$, respectively. Therefore $\Delta(E,F)\leq\eta$. ∎
7.2 LAN expansion and Gaussian equivalence
Fix $\pi\in(0,1)$ and let $k/n\to\pi$. Let $P_{n}=T_{n,k}$ and $Q_{n}=T_{n,k+1}$. Define the score direction $s=\Sigma_{\pi}^{+}v$ and the scalar score statistic
$Z_{n}:=s^{\top}\bigl(N-\mathbb{E}_{P_{n}}[N]\bigr).$
Then $Z_{n}$ is a sum of independent bounded increments (a triangular array) with $\mathrm{Var}(Z_{n})=nI_{\pi}+o(n)$.
Theorem 7.3 (LAN for the shuffle model).
Assume $\delta_{\mathrm{full}}>0$ and $k/n\to\pi\in(0,1)$. Let $\mu_{n}=\sqrt{I_{\pi}/n}$ and $\xi_{n}:=Z_{n}/\sqrt{nI_{\pi}}$. Then under $P_{n}$,
$\Lambda_{n,k}=\mu_{n}\xi_{n}-\frac{1}{2}\mu_{n}^{2}+o_{P_{n}}(1),\qquad\xi_{n}\Rightarrow N(0,1),$
and a Berry–Esseen bound holds: $\sup_{t}|\mathbb{P}(\xi_{n}\leq t)-\Phi(t)|\leq C/\sqrt{n}$.
Proof.
On the typical set $A_{n}$ from Theorem 4.1, $L_{n,k}=1+U_{n,k}$ with
$U_{n,k}=\frac{1}{n}Z_{n}+O(n^{-3/4}),$
uniformly on $A_{n}$, and $\mathbb{P}(A_{n}^{c})$ is exponentially small. Since $Z_{n}=O_{P_{n}}(\sqrt{n})$, we have $U_{n,k}=O_{P_{n}}(n^{-1/2})$ uniformly. Expand $\log(1+U_{n,k})=U_{n,k}-\frac{1}{2}U_{n,k}^{2}+O(U_{n,k}^{3})$, substitute $U_{n,k}=\mu_{n}\xi_{n}+o_{P_{n}}(1)$, and use $\mu_{n}^{2}=I_{\pi}/n\to 0$ to replace $\xi_{n}^{2}$ by $1$ in the quadratic term up to $o_{P_{n}}(1)$ (since $\xi_{n}\Rightarrow N(0,1)$ implies $\mu_{n}^{2}(\xi_{n}^{2}-1)\to^{P}0$ by Slutsky’s lemma). The Berry–Esseen bound for $\xi_{n}$ follows from a triangular-array Berry–Esseen inequality applied to the independent summands in $Z_{n}$ (bounded third moments from full support). ∎
Theorem 7.4 (Quantitative Le Cam distance to a Gaussian shift).
Let $E_{n}=(T_{n,k},T_{n,k+1})$ with $k/n\to\pi\in(0,1)$ and let $G_{\mu_{n}}=(N(0,1),N(\mu_{n},1))$ with $\mu_{n}=\sqrt{I_{\pi}/n}$. Then
$\Delta(E_{n},G_{\mu_{n}})\leq\frac{C}{\sqrt{n}}.$
Proof.
Theorem 6.4 yields $\sup_{\alpha}|\mathcal{T}_{E_{n}}(\alpha)-\mathcal{T}_{\mu_{n}}(\alpha)|\leq C/\sqrt{n}$. Apply Lemma 7.2. ∎
Corollary 7.5 (Asymptotic sufficiency of the score statistic).
Let $S_{n}$ be the experiment generated by the scalar statistic $Z_{n}$, i.e. $S_{n}=(\mathcal{L}_{T_{n,k}}(Z_{n}),\mathcal{L}_{T_{n,k+1}}(Z_{n}))$. Then $\Delta(E_{n},S_{n})\leq C/\sqrt{n}$. In particular, any test or privacy-relevant functional based on the full histogram can be matched (up to $O(n^{-1/2})$) by a procedure using only $Z_{n}$.
Proof.
By Theorem 7.3, $S_{n}$ is $O(n^{-1/2})$-close to the same Gaussian shift $G_{\mu_{n}}$. Combine with Theorem 7.4 and the triangle inequality. ∎
8 Exact finite-$n$ privacy curves and computable bounds
8.1 Exact positive-part formula
Proposition 8.1 (Exact positive-part formula).
Let $P,Q$ be distributions on a finite space with $Q\ll P$ and likelihood ratio $L=dQ/dP$. Then for every $\varepsilon\geq 0$,
$\delta_{Q\|P}(\varepsilon)=\sup_{A}\{Q(A)-e^{\varepsilon}P(A)\}=\mathbb{E}_{P}[(L-e^{\varepsilon})_{+}].$
The supremum is attained by the likelihood-ratio threshold event $A^{\star}=\{L>e^{\varepsilon}\}$ (ties can be randomized).
Proof.
For any $A$, $Q(A)-e^{\varepsilon}P(A)=\mathbb{E}_{P}[(L-e^{\varepsilon})\mathbf{1}_{A}]\leq\mathbb{E}_{P}[(L-e^{\varepsilon})_{+}]$. Taking $A=A^{\star}$ yields equality pointwise. ∎
Corollary 8.2 (Exact curve for the canonical shuffle pair).
Assume $\delta_{\star}>0$. For $(T_{n,0},T_{n,1})$ with $L_{n,0}=1+U_{n}$,
$\delta_{T_{n,1}\|T_{n,0}}(\varepsilon)=\mathbb{E}_{T_{n,0}}[(1+U_{n}-e^{\varepsilon})_{+}].$
8.2 Binary-output channels: one-dimensional binomial formulas
Assume $\mathcal{Y}=\{0,1\}$ with $p_{0}=W_{0}(1)\in(0,1)$ and $p_{1}=W_{1}(1)$. Under $T_{n,0}$ the histogram is determined by $K=N_{1}\sim\mathrm{Bin}(n,p_{0})$, and the likelihood ratio is affine in $K$:
$L_{n,0}(K)=\frac{n-K}{n}\frac{1-p_{1}}{1-p_{0}}+\frac{K}{n}\frac{p_{1}}{p_{0}}.$
Theorem 8.3 (Exact finite-$n$ privacy curve for binary-output channels).
For every $\varepsilon\geq 0$,
$\delta_{T_{n,1}\|T_{n,0}}(\varepsilon)=\sum_{k=0}^{n}\binom{n}{k}p_{0}^{k}(1-p_{0})^{n-k}\,\bigl(L_{n,0}(k)-e^{\varepsilon}\bigr)_{+}.$
Analogous one-dimensional formulas hold for $\mathrm{JSD}(T_{n,0}\|T_{n,1})$ and $D_{\alpha}(T_{n,1}\|T_{n,0})$ by substituting $D(L_{n,0}(k))$ and $L_{n,0}(k)^{\alpha}$.
Proof.
Apply Proposition 8.1 with $P=T_{n,0}$ and note that $L_{n,0}$ is a deterministic function of $K$. ∎
8.3 A computable Chernoff bound for general alphabets
Let $Z=r(Y)$ for $Y\sim W_{0}$ and define the mgf $M(\lambda)=\mathbb{E}[e^{\lambda Z}]$ and derivative $M^{\prime}(\lambda)=\mathbb{E}[Ze^{\lambda Z}]$. Let $\tau=e^{\varepsilon}-1$.
Theorem 8.4 (Chernoff-type bound for the one-sided curve).
Assume $\delta_{\star}>0$. For every $\lambda>0$,
$\delta_{T_{n,1}\|T_{n,0}}(\varepsilon)\leq\exp(-\lambda n\tau)\,M(\lambda)^{n-1}\,\bigl(M(\lambda)+M^{\prime}(\lambda)\bigr),$
and thus $\delta_{T_{n,1}\|T_{n,0}}(\varepsilon)$ is bounded by the infimum of the right-hand side over $\lambda>0$.
Proof.
By Corollary 8.2, $\delta_{T_{n,1}\|T_{n,0}}(\varepsilon)=\mathbb{E}[(U_{n}-\tau)_{+}]\leq\mathbb{E}[(1+U_{n})\mathbf{1}\{U_{n}\geq\tau\}]$. On $\{U_{n}\geq\tau\}$ we have $e^{\lambda n(U_{n}-\tau)}\geq 1$, hence $(1+U_{n})\mathbf{1}\{U_{n}\geq\tau\}\leq(1+U_{n})e^{\lambda n(U_{n}-\tau)}$. Writing $S_{n}=\sum_{i}Z_{i}=nU_{n}$ and using independence gives $\mathbb{E}[e^{\lambda S_{n}}]=M(\lambda)^{n}$ and $\mathbb{E}[S_{n}e^{\lambda S_{n}}]=nM^{\prime}(\lambda)M(\lambda)^{n-1}$, yielding the stated bound. ∎
9 Unbundled multi-message shuffling
We consider the unbundled model where each user sends $m\geq 1$ independent messages and the shuffler permutes all $nm$ messages individually.
9.1 Model
Fix $m\in\mathbb{N}$. In the unbundled $m$-message model, each user $i$ releases $m$ independent messages $Y_{i,1},\dots,Y_{i,m}\in\mathcal{Y}$ through the same local randomizer. The shuffler permutes the multiset of all $nm$ messages, and the analyst observes only the message-level histogram
$N^{(m)}=(N^{(m)}_{y})_{y\in\mathcal{Y}},\qquad\sum_{y\in\mathcal{Y}}N_{y}^{(m)}=nm.$
For a dataset with exactly $k$ ones, denote the law of $N^{(m)}$ by $T^{(m)}_{n,k}$, and write
$L^{(m)}_{n,k}:=\frac{dT^{(m)}_{n,k+1}}{dT^{(m)}_{n,k}},\qquad\Lambda^{(m)}_{n,k}:=\log L^{(m)}_{n,k}$.
9.2 Exact likelihood ratio for the canonical pair
Let $w(y)=W_{1}(y)/W_{0}(y)$ and consider the canonical pair $(T^{(m)}_{n,0},T^{(m)}_{n,1})$.
Theorem 9.1 (Exact likelihood ratio for unbundled $m$-message shuffling).
Let $N^{(m)}=(N_{y})$ be the released histogram of $nm$ messages. The likelihood ratio $L^{(m)}_{n,0}(N)$ equals the degree-$m$ symmetric polynomial
$L^{(m)}_{n,0}(N)=\frac{1}{\binom{nm}{m}}\sum_{\begin{subarray}{c}(m_{y})_{y\in\mathcal{Y}}:\\ 0\leq m_{y}\leq N_{y},\ \sum_{y}m_{y}=m\end{subarray}}\left(\prod_{y\in\mathcal{Y}}\binom{N_{y}}{m_{y}}\,w(y)^{m_{y}}\right)=\frac{[t^{m}]\prod_{y\in\mathcal{Y}}(1+w(y)t)^{N_{y}}}{\binom{nm}{m}}$,
where $t^{m}]$ extracts the coefficient of $t^{m}$. For $m=1$ this reduces to the linear identity $L^{(1)}_{n,0}(N)=\frac{1}{n}\sum_{y}N_{y}w(y)$.
Proof.
Under $T^{(m)}_{n,0}$ all $nm$ messages are i.i.d. from $W_{0}$, so the histogram pmf is multinomial. Under $T^{(m)}_{n,1}$, exactly $m$ messages (from the distinguished user) are i.i.d. from $W_{1}$ and the remaining nmmnm-m are i.i.d. from $W_{0}$, so the histogram is a convolution of two multinomials. Dividing by the null multinomial pmf cancels the $W_{0}$ factors and yields the stated coefficient-extraction identity. ∎
Theorem 9.2 (Explicit $(\varepsilon,\delta)$ bound for unbundled shuffling).
Assume $\delta_{\star}>0$ and let $w_{\max}=\max_{y\in\mathcal{Y}}w(y)$. For every $\varepsilon\geq 0$,
$\delta_{T^{(m)}_{n,1}\|T^{(m)}_{n,0}}(\varepsilon)\leq w_{\max}^{m}\exp\!\left(-\frac{2n(e^{\varepsilon}-1)^{2}}{w_{\max}^{2m}}\right).$
Proof.
By Proposition 8.1,
$\delta_{T^{(m)}_{n,1}\|T^{(m)}_{n,0}}(\varepsilon)\leq\mathbb{E}\bigl[L^{(m)}_{n,0}\mathbf{1}\{L^{(m)}_{n,0}\geq e^{\varepsilon}\}\bigr]\leq w_{\max}^{m}\,\mathbb{P}\bigl(L^{(m)}_{n,0}\geq e^{\varepsilon}\bigr),$
since $L^{(m)}_{n,0}\leq w_{\max}^{m}$ pointwise by Theorem 9.1. Under $T^{(m)}_{n,0}$, $L^{(m)}_{n,0}$ is a bounded order-$m$ U-statistic in the $nm$ i.i.d. message-level ratios $w(Y_{i})$ with mean 11, so Hoeffding’s inequality for bounded U-statistics (see [22, Section 5.6]), with range upper-bounded by $w_{\max}^{m}$ (since the kernel takes values in $w_{\min}^{m},w_{\max}^{m}]$ and $w_{\max}^{m}-w_{\min}^{m}\leq w_{\max}^{m}$), gives
$\mathbb{P}(L^{(m)}_{n,0}-1\geq t)\leq\exp\!\left(-\frac{2\lfloor nm/m\rfloor t^{2}}{w_{\max}^{2m}}\right)=\exp\!\left(-\frac{2nt^{2}}{w_{\max}^{2m}}\right).$
Set $t=e^{\varepsilon}-1$. ∎
9.3 Canonical unbundled GDP with Berry–Esseen rate
Theorem 9.3 (Unbundled multi-message GDP, canonical pair).
Assume full support $\delta_{\mathrm{full}}>0$ and fix $m\in\mathbb{N}$. Consider the canonical unbundled neighboring pair $(T^{(m)}_{n,0},T^{(m)}_{n,1})$. Let $\chi^{2}=\chi^{2}(W_{1}\|W_{0})$ and define
$\mu_{n}(m):=\sqrt{\frac{m\chi^{2}}{n}}.$
Then the privacy loss $\Lambda^{(m)}_{n,0}$ satisfies the LAN/GDP expansion
$\Lambda^{(m)}_{n,0}=\mu_{n}(m)\,\xi_{n}-\frac{1}{2}\mu_{n}(m)^{2}+o_{P}(1),\qquad\xi_{n}\Rightarrow N(0,1).$
Moreover, there exists $C>0$ depending only on $(d,\delta_{\mathrm{full}})$ and $m$ such that for all large enough $n$,
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{\Lambda^{(m)}_{n,0}+\mu_{n}(m)^{2}/2}{\mu_{n}(m)}\leq t\right)-\Phi(t)\right|\leq\frac{C}{\sqrt{n}}.$
Proof.
We give a complete proof via a U-statistic representation and Hoeffding decomposition.
Step 1: exact U-statistic representation.
Consider the ordered sample $(Y_{1},\dots,Y_{nm})$ of all $nm$ messages. Under $T^{(m)}_{n,0}$, all $nm$ messages are i.i.d. with law $W_{0}$. Under $T^{(m)}_{n,1}$, exactly $m$ of the $nm$ message positions correspond to the changed user and are drawn from $W_{1}$; the remaining (n1)m(n-1)m are drawn from $W_{0}$. Conditioning on the ordered sample under the null, the $m$ special positions are uniform among all $\binom{nm}{m}$ subsets, hence
$L^{(m)}_{n,0}(Y_{1},\dots,Y_{nm})=\frac{1}{\binom{nm}{m}}\sum_{1\leq i_{1}<\cdots<i_{m}\leq nm}\prod_{j=1}^{m}w(Y_{i_{j}}),$
a symmetric U-statistic of order $m$ with kernel $h(y_{1},\dots,y_{m})=\prod_{j=1}^{m}w(y_{j})$.
Step 2: Hoeffding decomposition.
Let $\theta=\mathbb{E}[h(Y_{1},\dots,Y_{m})]$ under $Y_{i}\sim W_{0}$ i.i.d. Since $\mathbb{E}[w(Y)]=\sum_{y}W_{1}(y)=1$, we have $\theta=1$. Define $V_{n}=L^{(m)}_{n,0}-1$. The first Hoeffding projection is
$h_{1}(y):=\mathbb{E}[h(y,Y_{2},\dots,Y_{m})]-\theta=w(y)-1=r(y).$
Therefore
$V_{n}=\frac{m}{nm}\sum_{i=1}^{nm}r(Y_{i})+R^{\mathrm{deg}}_{n}=\frac{1}{n}\sum_{i=1}^{nm}r(Y_{i})+R^{\mathrm{deg}}_{n},$
where $R^{\mathrm{deg}}_{n}$ is the sum of degenerate Hoeffding components of orders $2$ through $m$.
Step 3: variance scaling and negligibility of degenerate terms.
Let $X=r(Y)$ for $Y\sim W_{0}$, so $\mathbb{E}[X]=0$ and $\mathrm{Var}(X)=\chi^{2}$. The leading term equals $\frac{1}{n}\sum_{i=1}^{nm}X_{i}$, hence has variance
$\mathrm{Var}\!\left(\frac{1}{n}\sum_{i=1}^{nm}X_{i}\right)=\frac{m\chi^{2}}{n}=\mu_{n}(m)^{2}.$
For fixed $m$, $R^{\mathrm{deg}}_{n}$ is the sum of degenerate Hoeffding components of orders $2$ through $m$, computed over $nm$ observations. The dominant contribution (order $2$) has $\mathrm{Var}=O((nm)^{-2})$, and higher-order components have $\mathrm{Var}=O((nm)^{-j})$ for $j\geq 3$, so $\mathrm{Var}(R^{\mathrm{deg}}_{n})\leq C/(nm)^{2}$ (see [22, Chapter 5]), hence $R^{\mathrm{deg}}_{n}=O_{P}(n^{-1})$. Since $\mu_{n}(m)=\Theta(n^{-1/2})$, we have $R^{\mathrm{deg}}_{n}/\mu_{n}(m)\to 0$ in probability.
Step 4: Berry–Esseen for the leading projection.
Because $\delta_{\mathrm{full}}>0$ implies $r$ is bounded, $X$ has a bounded third absolute moment. Apply the classical Berry–Esseen theorem to $\sum_{i=1}^{nm}X_{i}$ to obtain Kolmogorov error $O(1/\sqrt{nm})=O(1/\sqrt{n})$ for the normalized leading term. Stability under the $o_{P}(1)$ perturbation $R^{\mathrm{deg}}_{n}/\mu_{n}(m)$ yields the same rate for $V_{n}/\mu_{n}(m)$.
Step 5: log transform.
Since $r$ is bounded and $m$ is fixed, $L^{(m)}_{n,0}$ is bounded in $L^{\infty}$ uniformly in $n$, so $(V_{n}^{2})$ is uniformly integrable. With $V_{n}=O_{P}(\mu_{n}(m))$ and $\mu_{n}(m)\to 0$,
$\Lambda^{(m)}_{n,0}=\log(1+V_{n})=V_{n}-\frac{1}{2}V_{n}^{2}+o_{P}(\mu_{n}(m)^{2}).$
Taking expectations yields $\mathbb{E}[\Lambda^{(m)}_{n,0}]=-\frac{1}{2}\mu_{n}(m)^{2}+o(\mu_{n}(m)^{2})$, and dividing by $\mu_{n}(m)$ gives the LAN/GDP expansion. The Kolmogorov bound transfers from $V_{n}/\mu_{n}(m)$ to the centered-and-scaled $\Lambda^{(m)}_{n,0}$ by monotonicity of $\log$ and the same stability argument. ∎
9.4 Proportional compositions: unbundled GDP (fixed $m$)
We now extend to $k/n\to\pi\in(0,1)$ with fixed $m$.
Theorem 9.4 (Unbundled multi-message GDP, proportional compositions).
Assume full support and fix $m\in\mathbb{N}$. Let $k=k(n)$ satisfy $k/n\to\pi\in(0,1)$. Then the unbundled neighboring pair $(T^{(m)}_{n,k},T^{(m)}_{n,k+1})$ is asymptotically GDP with parameter
$\mu_{n,\pi}(m):=\sqrt{\frac{mI_{\pi}}{n}},\qquad I_{\pi}=v^{\top}\Sigma_{\pi}^{+}v,$
with $\Sigma_{\pi}$ the fixed-composition covariance (1). More precisely,
$\Lambda^{(m)}_{n,k}=\mu_{n,\pi}(m)\,\xi_{n}-\frac{1}{2}\mu_{n,\pi}(m)^{2}+o_{P}(1),\qquad\xi_{n}\Rightarrow N(0,1),$
and the Kolmogorov distance to the $N(-\mu^{2}/2,\mu^{2})$ limit is $O(n^{-1/2})$ for fixed $m$.
Proof.
We keep the structure parallel to the canonical case, but the leading projection is now a fixed-composition score.
Step 1: conditional-expectation representation at the message level.
Under $T^{(m)}_{n,k}$, select the distinguished user whose bit flips when moving to $k+1$. Let $N_{\star}\in\mathbb{Z}_{\geq 0}^{d}$ be the histogram of the $m$ messages produced by this user, and let $S$ be the histogram of the remaining (n1)m(n-1)m messages. Then
$N^{(m)}=N_{\star}+S,$
and $S$ has the same law under $T^{(m)}_{n,k}$ and $T^{(m)}_{n,k+1}$. Moreover,
$N_{\star}\sim\mathrm{Mult}(m,W_{0})\ \text{under}\ T^{(m)}_{n,k},\qquad N_{\star}\sim\mathrm{Mult}(m,W_{1})\ \text{under}\ T^{(m)}_{n,k+1}$,
and $N_{\star}$ is independent of $S$. Therefore
$L^{(m)}_{n,k}(N^{(m)})=\mathbb{E}\!\left[\left.\prod_{y\in\mathcal{Y}}w(y)^{(N_{\star})_{y}}\ \right|\ N^{(m)}\right]=\mathbb{E}\!\left[\left.\prod_{j=1}^{m}w(Y^{\star}_{j})\ \right|\ N^{(m)}\right],$ (25)
where $Y^{\star}_{1},\dots,Y^{\star}_{m}$ are the distinguished messages.
Expand $\prod_{j=1}^{m}(1+r(Y^{\star}_{j}))$:
$\prod_{j=1}^{m}(1+r(Y^{\star}_{j}))=1+\sum_{j=1}^{m}r(Y^{\star}_{j})+R_{\star},$
where $R_{\star}$ is a finite sum of monomials of degree at least 22 in the bounded variables $r(Y^{\star}_{j})$. Taking conditional expectations in (25),
$L^{(m)}_{n,k}-1=\sum_{j=1}^{m}\mathbb{E}[r(Y^{\star}_{j})\mid N^{(m)}]+\mathbb{E}[R_{\star}\mid N^{(m)}].$ (26)
By symmetry of the $m$ distinguished messages, $\mathbb{E}[r(Y^{\star}_{j})\mid N^{(m)}]$ is the same for each $j$. With $\pi_{n}=k/(n-1)\to\pi$ as in Section 4, let $s_{\pi_{n}}:=\Sigma_{\pi_{n}}^{+}v$.
Step 2: posterior tilt and linear term.
We first establish a posterior tilt approximation, then extract the linear term.
Posterior representation. By Bayes’ rule and $N^{(m)}=N_{\star}+S$, with $\mathcal{N}$ denoting a fixed histogram with $\|\mathcal{N}\|_{1}=nm$,
$\mathbb{P}(N_{\star}=a\mid N^{(m)}=\mathcal{N})\propto\mathrm{Mult}(m,W_{0})(a)\cdot\mathbb{P}(S=\mathcal{N}-a),$
over $a\in\mathbb{Z}_{\geq 0}^{d}$ with $\|a\|_{1}=m$ and $0\leq a\leq\mathcal{N}$.
Reference histogram and iterated ratio lemma. Fix a reference histogram $a_{0}\in\mathbb{Z}_{\geq 0}^{d}$ with $\|a_{0}\|_{1}=m$ (for instance the nearest integer rounding of $mW_{0}$ adjusted to have sum $m$). Both $\mathcal{N}-a$ and $\mathcal{N}-a_{0}$ have $\ell^{1}$-norm (n1)m(n-1)m, matching $\|S\|_{1}$. For any admissible $a$, the difference $a_{0}-a$ satisfies $\mathbf{1}^{\top}(a_{0}-a)=0$ and $\|a_{0}-a\|_{1}\leq 2m$. Decompose $a_{0}-a$ into at most $m$ elementary transpositions $t_{1},\dots,t_{\ell}$ ($\ell\leq m$), each with $\|t_{j}\|_{1}=2$ and $\mathbf{1}^{\top}t_{j}=0$. Apply Lemma A.2 (Appendix A) iteratively at sample size (n1)m(n-1)m: on the typical set
$\mathcal{T}_{n,M}:=\left\{\|(\mathcal{N}-\mathbb{E}[N^{(m)}])^{-}\|_{2}\leq M\sqrt{n}\right\},$
(here the $\ell_{2}$-norm is used for compatibility with the Edgeworth ratio lemma; the probability $\mathbb{P}(\mathcal{T}_{n,M}^{c})$ is controlled by tightness of $(N^{(m)}-\mathbb{E}[N^{(m)}])/\sqrt{n}$, see below) the improved remainder of Lemma A.2 gives $O(1/n)$ per step (since the standardized deviation is $O(\sqrt{n}/\sqrt{nm})=O(1/\sqrt{m})=O(1)$), and each intermediate point remains in the window because $\|t_{j}\|_{1}=2=O(1)$. Summing over $\ell\leq m$ steps and using that $m$ is fixed,
$\log\frac{\mathbb{P}(S=\mathcal{N}-a)}{\mathbb{P}(S=\mathcal{N}-a_{0})}=-\frac{1}{(n-1)m}(a_{0}-a)^{\top}\Sigma_{\pi_{n}}^{+}\bigl((\mathcal{N}-a_{0})-\mathbb{E}[S]\bigr)+O(1/n),$ (27)
uniformly over all admissible $a$ on $\mathcal{T}_{n,M}$.
Exponential tilt approximation. Define the tilt direction
$\theta:=\frac{1}{(n-1)m}\,\Sigma_{\pi_{n}}^{+}\bigl(\mathcal{N}-\mathbb{E}[N^{(m)}]\bigr),$
which satisfies $\|\theta\|=O(n^{-1/2})$ on $\mathcal{T}_{n,M}$. Since $(\mathcal{N}-a_{0})-\mathbb{E}[S]=(\mathcal{N}-\mathbb{E}[N^{(m)}])+(mW_{0}-a_{0})$ and $\Sigma_{\pi_{n}}^{+}(mW_{0}-a_{0})$ is a bounded constant, (27) becomes
$\frac{\mathbb{P}(S=\mathcal{N}-a)}{\mathbb{P}(S=\mathcal{N}-a_{0})}=\exp\bigl(\theta^{\top}(a-a_{0})+O(1/n)\bigr)=\exp\bigl(\theta^{\top}a+O(1/n)\bigr)\cdot e^{-\theta^{\top}a_{0}}.$
Let $W_{0}^{\theta}(y)\propto W_{0}(y)e^{\theta_{y}}$ be the exponential tilt. Since the number of histograms with $\|a\|_{1}=m$ is $\binom{m+d-1}{d-1}$ (a constant for fixed $m,d$), the uniform $O(1/n)$ bound implies
$\left\|\mathcal{L}(N_{\star}\mid N^{(m)}=\mathcal{N})-\mathrm{Mult}(m,W_{0}^{\theta})\right\|_{\mathrm{TV}}\leq\frac{C^{\prime}}{n}$ (28)
on $\mathcal{T}_{n,M}$, for a constant $C^{\prime}=C^{\prime}(m,d,\delta_{\mathrm{full}})$. Lifting from the histogram $N_{\star}$ to the ordered tuple $(Y^{\star}_{1},\dots,Y^{\star}_{m})$ multiplies total variation by at most the maximal fiber size $m!$, hence
$\left\|\mathcal{L}\bigl((Y^{\star}_{1},\dots,Y^{\star}_{m})\mid N^{(m)}=\mathcal{N}\bigr)-\bigotimes_{j=1}^{m}W_{0}^{\theta}\right\|_{\mathrm{TV}}\leq\frac{m!\,C^{\prime}}{n}.$
Since $(N^{(m)}-\mathbb{E}[N^{(m)}])/\sqrt{n}$ is tight, for every $\eta>0$ one can choose $M$ so that $\mathbb{P}(\mathcal{T}_{n,M})\geq 1-\eta$ for all large $n$.
Linearization of each conditional expectation. Under the tilt $W_{0}^{\theta}$, since $\mathbb{E}[r(Y)]=0$ under $W_{0}$ (because $\sum_{y}r(y)W_{0}(y)=\sum_{y}v(y)=0$), a first-order expansion gives
$\mathbb{E}_{W_{0}^{\theta}}[r(Y)]=\sum_{y}r(y)W_{0}(y)\theta_{y}+O(\|\theta\|^{2})=v^{\top}\theta+O(\|\theta\|^{2}),$
where we used $r(y)W_{0}(y)=v(y)$. Combining with the TV bound (28),
$\mathbb{E}[r(Y^{\star}_{1})\mid N^{(m)}]=v^{\top}\theta+O_{P}(n^{-1}).$
Substituting $\theta=(1/((n-1)m))\Sigma_{\pi_{n}}^{+}(\mathcal{N}-\mathbb{E}[N^{(m)}])$ yields $v^{\top}\theta=(1/((n-1)m))s_{\pi_{n}}^{\top}(\mathcal{N}-\mathbb{E}[N^{(m)}])$ with $s_{\pi_{n}}=\Sigma_{\pi_{n}}^{+}v$. Replacing $(n-1)^{-1}$ by $n^{-1}$ introduces an error $O(n^{-2})\cdot O_{P}(\sqrt{n})=O_{P}(n^{-3/2})$, so
$\mathbb{E}[r(Y^{\star}_{1})\mid N^{(m)}]=\frac{1}{nm}\,s_{\pi_{n}}^{\top}\Bigl(N^{(m)}-\mathbb{E}[N^{(m)}]\Bigr)+O_{P}(n^{-1}),$ (29)
uniformly on $\mathcal{T}_{n,M}$ (constants depend on $d,\delta_{\mathrm{full}},m$). Multiplying by $m$ (using symmetry of the $m$ distinguished messages) yields
$\sum_{j=1}^{m}\mathbb{E}[r(Y^{\star}_{j})\mid N^{(m)}]=\frac{1}{n}\,s_{\pi_{n}}^{\top}\Bigl(N^{(m)}-\mathbb{E}[N^{(m)}]\Bigr)+O_{P}(n^{-1}).$ (30)
Step 3: control of the quadratic-and-higher remainder.
We claim
$\mathbb{E}[R_{\star}\mid N^{(m)}]=O_{P}(n^{-1}).$ (31)
Since $m$ is fixed, it suffices to show that for each fixed pair $j\neq\ell$, $\mathbb{E}[r(Y^{\star}_{j})r(Y^{\star}_{\ell})\mid N^{(m)}]=O_{P}(n^{-1})$, and similarly for higher-order monomials. Under the product tilt from Step 2, the $Y^{\star}_{j}$ are independent and $\mathbb{E}[r(Y^{\star}_{j})\mid N^{(m)}]=O_{P}(n^{-1/2})$ (the linearization scale), while $\mathbb{E}[r(Y^{\star}_{j})^{2}\mid N^{(m)}]=O(1)$ by boundedness. Thus
$\mathbb{E}[r(Y^{\star}_{j})r(Y^{\star}_{\ell})\mid N^{(m)}]=\mathbb{E}[r(Y^{\star}_{j})\mid N^{(m)}]\mathbb{E}[r(Y^{\star}_{\ell})\mid N^{(m)}]+O_{P}(n^{-1})=O_{P}(n^{-1}).$
The $O_{P}(n^{-1})$ error absorbs both the independence approximation error ($O(1/n)$ in total variation times bounded integrand) and the product-of-means term ($O_{P}(n^{-1})$). The same reasoning applies to higher-order monomials (they contain at least two factors with $O_{P}(n^{-1/2})$ means), proving (31).
Step 4: asymptotic normality and variance identification.
Combine (26), (30), and (31):
$V_{n}:=L^{(m)}_{n,k}-1=\frac{1}{n}\,s_{\pi_{n}}^{\top}\Bigl(N^{(m)}-\mathbb{E}[N^{(m)}]\Bigr)+O_{P}(n^{-1}).$ (32)
Let $H_{n}:=s_{\pi_{n}}^{\top}(N^{(m)}-\mathbb{E}[N^{(m)}])$. By a multivariate CLT for fixed-alphabet histograms under fixed composition,
$\frac{H_{n}}{\sqrt{nm}}\Rightarrow N(0,s^{\top}\Sigma_{\pi}s)=N(0,I_{\pi}),$
with $s=\Sigma_{\pi}^{+}v$. Moreover $\mathrm{Var}(H_{n})=nm\,I_{\pi}+o(n)$ and a one-dimensional Berry–Esseen bound of order $O(n^{-1/2})$ holds for $H_{n}$ (bounded third moments, fixed $m$). Thus
$\mathrm{Var}(V_{n})=\frac{1}{n^{2}}\mathrm{Var}(H_{n})+o(n^{-1})=\frac{mI_{\pi}}{n}+o(n^{-1}),$
matching $\mu_{n,\pi}(m)^{2}$.
Step 5: log expansion and GDP limit.
Since $r$ is bounded and $m$ fixed, $L^{(m)}_{n,k}$ is bounded in $L^{\infty}$ uniformly in $n$, hence $(V_{n}^{2})$ is uniformly integrable. With $V_{n}=O_{P}(\mu_{n,\pi}(m))$ and $\mu_{n,\pi}(m)\to 0$,
$\Lambda^{(m)}_{n,k}=\log(1+V_{n})=V_{n}-\frac{1}{2}V_{n}^{2}+o_{P}(\mu_{n,\pi}(m)^{2}).$
The centered-and-scaled $\Lambda^{(m)}_{n,k}$ therefore inherits asymptotic normality and the $O(n^{-1/2})$ Kolmogorov bound from $V_{n}/\mu_{n,\pi}(m)$ by the same argument as in Theorem 9.3. ∎
9.5 Bundled vs. unbundled: a quantitative comparison
In the bundled model, the $m$-tuple is treated as a single message in alphabet $\mathcal{Y}^{m}$ and shuffling permutes $n$ bundled outputs. These two models have different privacy behavior; unbundling is strictly better for fixed $m\geq 2$.
Lemma 9.5 (Chi-square multiplicativity for product channels).
Let $W_{0}^{\otimes m}$ and $W_{1}^{\otimes m}$ be the product distributions on $\mathcal{Y}^{m}$. Then
$1+\chi^{2}(W_{1}^{\otimes m}\|W_{0}^{\otimes m})=\bigl(1+\chi^{2}(W_{1}\|W_{0})\bigr)^{m}.$
Proof.
Let $Y_{1},\dots,Y_{m}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}W_{0}$ and set $w(y)=W_{1}(y)/W_{0}(y)$. The product-channel likelihood ratio is $\prod_{j=1}^{m}w(Y_{j})$, so
$\chi^{2}(W_{1}^{\otimes m}\|W_{0}^{\otimes m})=\mathbb{E}\Bigl[\Bigl(\prod_{j=1}^{m}w(Y_{j})-1\Bigr)^{2}\Bigr]=\mathbb{E}\Bigl[\prod_{j=1}^{m}w(Y_{j})^{2}\Bigr]-1=\bigl(\mathbb{E}[w(Y)^{2}]\bigr)^{m}-1,$
and $\mathbb{E}[w(Y)^{2}]=1+\chi^{2}(W_{1}\|W_{0})$ by definition. ∎
Proposition 9.6 (Unbundled is strictly better than bundled for $m\geq 2$).
Assume $\chi^{2}(W_{1}\|W_{0})>0$ and fix $m\geq 2$. In the bundled model (treating $\mathcal{Y}^{m}$ as the alphabet and shuffling $n$ bundled messages), the canonical GDP parameter satisfies
$\mu_{\mathrm{bund}}(m)^{2}=\frac{\chi^{2}(W_{1}^{\otimes m}\|W_{0}^{\otimes m})}{n}=\frac{(1+\chi^{2})^{m}-1}{n}.$
In the unbundled model, Theorem 9.3 gives
$\mu_{\mathrm{unb}}(m)^{2}=\frac{m\chi^{2}}{n}.$
For every $\chi^{2}>0$ and $m\geq 2$,
$\mu_{\mathrm{unb}}(m)^{2}<\mu_{\mathrm{bund}}(m)^{2},\qquad\text{equivalently}\qquad m\chi^{2}<(1+\chi^{2})^{m}-1.$
Since a smaller GDP parameter $\mu$ corresponds to stronger privacy (a tighter $(\varepsilon,\delta)$ curve), unbundled shuffling provides strictly better privacy than bundled shuffling for $m\geq 2$.
Proof.
The formulas are Lemma 9.5 and Theorem 9.3. For the strict inequality, expand $(1+\chi^{2})^{m}-1=m\chi^{2}+\sum_{j=2}^{m}\binom{m}{j}(\chi^{2})^{j}$ and note that the sum over $j\geq 2$ is strictly positive. In particular, the ratio of bundled to unbundled GDP parameter squared satisfies
$\frac{\mu_{\mathrm{bund}}^{2}}{\mu_{\mathrm{unb}}^{2}}=\frac{(1+\chi^{2})^{m}-1}{m\chi^{2}}\geq 1+\frac{(m-1)\chi^{2}}{2},$
giving practitioners a quick lower bound on the privacy gap. ∎
10 The Gaussian regime boundary (randomized response)
We quantify how Berry–Esseen accuracy deteriorates as a local privacy parameter grows with $n$. This provides an analytic bridge between the sub-critical Gaussian regime treated above and the critical Poisson/Skellam limits in the companion paper (Part II).
10.1 Setup: randomized response and the boundary parameter
Specialize to binary randomized response (RR) [29] with local privacy parameter $\varepsilon_{0}=\varepsilon_{0}(n)\geq 0$ and output alphabet $\mathcal{Y}=\{0,1\}$. Under input 0, RR outputs $1$ with probability
$q_{n}:=\frac{1}{1+e^{\varepsilon_{0}}}\in(0,1/2],$
and 0 otherwise; under input $1$ the probabilities are swapped. Thus
$W_{0}(1)=q_{n},\quad W_{0}(0)=1-q_{n},\qquad W_{1}(1)=1-q_{n},\quad W_{1}(0)=q_{n},$
and the pointwise likelihood ratios are $w(1)=e^{\varepsilon_{0}}$ and $w(0)=e^{-\varepsilon_{0}}$. The boundary parameter is
$a_{n}:=\frac{e^{\varepsilon_{0}}}{n}.$
The Gaussian regime corresponds to $a_{n}\to 0$, the critical regime to $a_{n}\to c^{2}\in(0,\infty)$, and the super-critical regime to $a_{n}\to\infty$.
10.2 Boundary Berry–Esseen
Theorem 10.1 (Boundary Berry–Esseen correction for RR).
Consider the single-message canonical neighboring pair $(T_{n,0},T_{n,1})$ in binary RR with local privacy parameter $\varepsilon_{0}=\varepsilon_{0}(n)$ and boundary scaling $a_{n}=e^{\varepsilon_{0}}/n$. Let $\Lambda_{n}=\log L_{n,0}(N)$ and $\mu_{n}=\sqrt{\chi^{2}/n}$ with $\chi^{2}=\chi^{2}(W_{1}\|W_{0})$. If $a_{n}\to 0$, then there exists a universal constant $C>0$ such that for all sufficiently large $n$,
$\sup_{t\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{\Lambda_{n}+\mu_{n}^{2}/2}{\mu_{n}}\leq t\right)-\Phi(t)\right|\leq C\sqrt{a_{n}}.$
Proof.
We follow a five-step argument and track the dependence on $a_{n}$.
Step 1: moments of the RR increment.
Let $X=r(Y)$ under $Y\sim W_{0}$. Then $X$ takes two values
$x_{+}:=e^{\varepsilon_{0}}-1,\qquad x_{-}:=e^{-\varepsilon_{0}}-1=-\frac{x_{+}}{e^{\varepsilon_{0}}},$
with probabilities $q_{n}$ and $1-q_{n}$. One checks $\mathbb{E}[X]=0$. The variance is
$\sigma^{2}:=\mathrm{Var}(X)=\mathbb{E}[X^{2}]=\frac{(e^{\varepsilon_{0}}-1)^{2}}{e^{\varepsilon_{0}}},$
so $\chi^{2}=\sigma^{2}$ and $\mu_{n}=\sigma/\sqrt{n}$. The third absolute moment is
$\rho_{3}:=\mathbb{E}|X|^{3}=(e^{\varepsilon_{0}}-1)^{3}\cdot\frac{1+e^{-2\varepsilon_{0}}}{1+e^{\varepsilon_{0}}}.$
Step 2: Lyapunov ratio and its $\sqrt{a_{n}}$ scaling.
A direct computation yields
$\frac{\rho_{3}}{\sigma^{3}}=e^{3\varepsilon_{0}/2}\cdot\frac{1+e^{-2\varepsilon_{0}}}{1+e^{\varepsilon_{0}}}\leq 2e^{\varepsilon_{0}/2},$
since $(1+e^{-2\varepsilon_{0}})\leq 2$ and $(1+e^{\varepsilon_{0}})\geq e^{\varepsilon_{0}}$. Therefore
$\frac{\rho_{3}}{\sigma^{3}\sqrt{n}}\leq 2\frac{e^{\varepsilon_{0}/2}}{\sqrt{n}}=2\sqrt{\frac{e^{\varepsilon_{0}}}{n}}=2\sqrt{a_{n}}.$
Step 3: Berry–Esseen for $U_{n}=n^{-1}\sum_{i=1}^{n}X_{i}$.
In the canonical RR shuffle pair, the exact identity $L_{n,0}=1+U_{n}$ holds (Lemma 3.2). Let $Z_{n}=\sum_{i=1}^{n}X_{i}/(\sigma\sqrt{n})=U_{n}/\mu_{n}$. By the classical Berry–Esseen theorem,
$\sup_{t}|\mathbb{P}(Z_{n}\leq t)-\Phi(t)|\leq C_{\mathrm{BE}}\frac{\rho_{3}}{\sigma^{3}\sqrt{n}}\leq 2C_{\mathrm{BE}}\sqrt{a_{n}}.$
Step 4: log transform preserves the $\sqrt{a_{n}}$ rate.
Let
$T_{n}:=\frac{\Lambda_{n}+\mu_{n}^{2}/2}{\mu_{n}}=\frac{\log(1+\mu_{n}Z_{n})+\mu_{n}^{2}/2}{\mu_{n}}=:g_{\mu_{n}}(Z_{n}),$
where $g_{\mu}(z)=\mu^{-1}\log(1+\mu z)+\mu/2$ is increasing. For any $t$,
$\mathbb{P}(T_{n}\leq t)=\mathbb{P}(Z_{n}\leq g_{\mu_{n}}^{-1}(t)),\qquad g_{\mu}^{-1}(t)=\frac{e^{\mu t-\mu^{2}/2}-1}{\mu}.$
Decompose
$|\mathbb{P}(T_{n}\leq t)-\Phi(t)|\leq|\mathbb{P}(Z_{n}\leq g_{\mu_{n}}^{-1}(t))-\Phi(g_{\mu_{n}}^{-1}(t))|+|\Phi(g_{\mu_{n}}^{-1}(t))-\Phi(t)|.$
The first term is bounded by the Kolmogorov distance for $Z_{n}$, hence $O(\sqrt{a_{n}})$. For the second term, a Taylor expansion shows $\sup_{t}|\Phi(g_{\mu}^{-1}(t))-\Phi(t)|\leq C\mu$ for a universal $C$, and for RR one has $\mu_{n}\leq e^{\varepsilon_{0}/2}/\sqrt{n}=\sqrt{a_{n}}$. Thus the second term is also $O(\sqrt{a_{n}})$.
Step 5: conclusion.
Combine the two bounds and absorb constants into a universal $C$. ∎
Remark 10.2 (Three regimes for shuffled randomized response).
Three regimes arise in terms of the boundary parameter $a_{n}=e^{\varepsilon_{0}}/n$: sub-critical ($a_{n}\to 0$, Gaussian limits and GDP amplification), critical ($a_{n}\to c^{2}\in(0,\infty)$, Poisson/Skellam limits), and super-critical ($a_{n}\to\infty$, separation with $\mathrm{TV}\to 1$). The latter two regimes are treated in the companion paper [23].
11 Application: frequency estimation
We include a practitioner-facing application showing that exact privacy calibration yields less noise than a general amplification bound for shuffled randomized response.
11.1 Estimator and error
Consider binary frequency estimation: each user holds $x_{i}\in\{0,1\}$ and we wish to estimate $p=n^{-1}\sum_{i}x_{i}$. With RR local privacy parameter $\varepsilon_{0}>0$, define the flip probability $q=(1+e^{\varepsilon_{0}})^{-1}$ and output $Y_{i}\in\{0,1\}$ by flipping with probability $q$. The analyzer observes $K=\sum_{i}\mathbf{1}\{Y_{i}=1\}$. An unbiased estimator is
$\hat{p}=\frac{K/n-q}{1-2q}.$
Since $\mathrm{Var}(K/n)\leq 1/(4n)$ for any $p$, the worst-case MSE satisfies
$\sup_{p\in[0,1]}\mathbb{E}[(\hat{p}-p)^{2}]\leq\frac{1}{4n(1-2q)^{2}},$
which improves monotonically with larger $\varepsilon_{0}$ (smaller $q$).
11.2 Exact privacy calibration for shuffled RR
For RR the shuffle likelihood ratio for the canonical neighboring pair is affine:
$L_{n,0}(K)=e^{-\varepsilon_{0}}+\frac{K}{n}\bigl(e^{\varepsilon_{0}}-e^{-\varepsilon_{0}}\bigr),$
and the tight two-sided curve $\delta(\varepsilon)$ can be computed exactly using the binomial formulas in Theorem 8.3 (and its two-sided variant). In contrast, amplification upper bounds such as Feldman–McMillan–Talwar [13] provide conservative calibrations that do not use the exact likelihood-ratio structure.
12 Discussion and open problems
Gaussian universality summary.
For fixed finite-output local randomizers with full support, the neighboring shuffle experiment is governed by a one-dimensional Gaussian shift with signal-to-noise parameter
$\mu_{n}(\pi)=\sqrt{\frac{I_{\pi}}{n}},\qquad I_{\pi}=v^{\top}\Sigma_{\pi}^{+}v,$
and, in the unbundled $m$-message model, by $\mu_{n,\pi}(m)=\sqrt{mI_{\pi}/n}$ for fixed $m$. This experiment-level statement explains the universality of leading constants for smooth divergences, yields sharp Jensen–Shannon and GDP constants, and provides an explicit limiting $(\varepsilon,\delta)$ curve with quantitative $O(n^{-1/2})$ control in the full-support regime.
How to use the constants in practice.
Given $(W_{0},W_{1})$, a sample size $n$, and a composition proportion $\pi\in(0,1)$, one first computes $v=W_{1}-W_{0}$ and the within-type covariances $\Sigma_{b}=\mathrm{diag}(W_{b})-W_{b}W_{b}^{\top}$. Next form the fixed-composition covariance $\Sigma_{\pi}=(1-\pi)\Sigma_{0}+\pi\Sigma_{1}$ and the Fisher constant $I_{\pi}=v^{\top}\Sigma_{\pi}^{+}v$. The corresponding GDP parameter is $\mu=\sqrt{I_{\pi}/n}$ (and in the canonical case $\mu=\sqrt{\chi^{2}(W_{1}\|W_{0})/n}$); in the unbundled multi-message setting, the parameter multiplies by $\sqrt{m}$ when $m$ is fixed. One may then convert $\mu$ to an $(\varepsilon,\delta)$ curve using the closed-form GDP trade-off (24). For finite-$n$ calibration, the exact positive-part identity in Proposition 8.1 reduces the one-sided curve $\delta_{Q\|P}(\varepsilon)$ to an expectation under the null (and the two-sided curve follows by taking the maximum over both directions), and for binary-output channels Theorem 8.3 provides an explicit one-dimensional binomial sum. Numerical comparisons confirming the quantitative advantage of exact GDP calibration over amplification bounds will appear in a companion numerical study.
Open problems.
A first problem in earlier versions was to obtain a uniform $O(n^{-2})$ Jensen–Shannon remainder for proportional compositions $k=\Theta(n)$. This is resolved in Theorem 4.5. A remaining challenge is to develop explicit computable $n^{-2}$ correction terms (and analogous corrections for general $f$-divergences) in terms of Edgeworth polynomials of the score statistic, which would improve finite-$n$ guidance beyond the leading Fisher constant.
A second problem is to treat growing alphabets $d=d_{n}\to\infty$. In that regime, lattice local CLT and Edgeworth expansions must be made dimension-dependent, and new growth conditions are needed to balance combinatorial factors against the $n^{-1/2}$ Gaussian approximation rates.
A third problem is to develop systematic Edgeworth corrections to the LAN and GDP limits. Second-order expansions for the privacy loss process would yield simultaneous $O(n^{-1})$ corrections for trade-off functions and privacy curves, improving finite-$n$ guidance across the Gaussian regime.
Pointer to non-Gaussian regimes.
Critical and super-critical regimes, where the local privacy parameter grows with $n$ and Gaussian limits can fail, are treated in the companion paper [23].
Appendix A Edgeworth ratio lemma
This appendix supplies the technical ratio lemma invoked in the proof of Theorem 4.1. The key input is a first-order lattice Edgeworth expansion for sums of independent one-hot vectors, specialized to the simplex lattice $\{x\in\mathbb{Z}^{d}:\mathbf{1}^{\top}x=m\}$ arising from histograms.
A.1 A local CLT with first-order Edgeworth correction
Let $d\geq 2$ be fixed and let $X_{1},\dots,X_{m}$ be independent random vectors taking values in $\{e_{1},\dots,e_{d}\}\subset\mathbb{R}^{d}$ (one-hot vectors). Write
$S_{m}:=\sum_{i=1}^{m}X_{i}\in\mathbb{Z}^{d},\qquad\mu:=\mathbb{E}[S_{m}],\qquad\Sigma_{\mathrm{ps}}:=\frac{1}{m}\mathrm{Cov}(S_{m}).$
We refer to $\Sigma_{\mathrm{ps}}$ as the per-sample covariance, so that $\mathrm{Cov}(S_{m})=m\Sigma_{\mathrm{ps}}$. In the fixed-composition setting of the main text, $\Sigma_{\mathrm{ps}}$ equals the fixed-composition covariance
$\Sigma_{\mathrm{ps}}=(1-\pi)\Sigma_{0}+\pi\Sigma_{1},$ (33)
consistent with (1). Since $\mathbf{1}^{\top}S_{m}=m$ almost surely, hence $\Sigma_{\mathrm{ps}}$ is singular with $\ker(\Sigma_{\mathrm{ps}})=\mathrm{span}\{\mathbf{1}\}$ and has rank d1d-1 under full support. To work with a nonsingular covariance, drop the last coordinate: for $x\in\mathbb{R}^{d}$, set $x^{-}=(x_{1},\dots,x_{d-1})\in\mathbb{R}^{d-1}$ and define the per-sample minor
$\Sigma_{\mathrm{ps}}^{-}:=\frac{1}{m}\mathrm{Cov}\bigl((S_{m})^{-}\bigr)\in\mathbb{R}^{(d-1)\times(d-1)}.$
Under full support, $\Sigma_{\mathrm{ps}}^{-}$ is positive definite. We impose a uniform full-support condition: there exists $p>0$ such that for all $i\leq m$ and all $y\in[d]$,
$\mathbb{P}(X_{i}=e_{y})\geq p.$ (34)
Lemma A.1 (Local CLT with first-order Edgeworth correction).
Assume (34). There exist constants $C,c>0$ (depending only on $d$ and $p$) and, for each $m$, a cubic polynomial $Q_{m}:\mathbb{R}^{d-1}\to\mathbb{R}$ with coefficients bounded uniformly in $m$, such that the following holds. For every $s\in\mathbb{Z}^{d}$ with $\mathbf{1}^{\top}s=m$ and
$z:=\frac{s-\mu}{\sqrt{m}},\qquad z^{-}:=\frac{s^{-}-\mu^{-}}{\sqrt{m}},$
whenever $\|z^{-}\|_{2}\leq m^{1/8}$ we have the Edgeworth local expansion
$\mathbb{P}(S_{m}=s)=\frac{1}{(2\pi m)^{(d-1)/2}\sqrt{\det(\Sigma_{\mathrm{ps}}^{-})}}\,\exp\!\left(-\frac{1}{2}(z^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}z^{-}\right)\left(1+m^{-1/2}Q_{m}(z^{-})+R_{m}(z^{-})\right),$ (35)
with remainder bound
$|R_{m}(z^{-})|\leq Cm^{-1}\bigl(1+\|z^{-}\|_{2}^{6}\bigr),\qquad|Q_{m}(z^{-})|\leq C\bigl(1+\|z^{-}\|_{2}^{3}\bigr).$
Proof.
This is a lattice Edgeworth expansion (local CLT with a polynomial correction) for sums of independent lattice-valued random vectors with bounded third moments and a uniform aperiodicity condition. In our categorical setting, all moments are bounded (each $X_{i}$ takes values in $\{e_{1},\dots,e_{d}\}$), uniform nondegeneracy $\Sigma_{\mathrm{ps}}^{-}\succ 0$ follows from (34) (every category has probability $\geq p$), and strong aperiodicity on the dual lattice holds because (34) ensures positive mass in all $d$ directions, which prevents the characteristic function from approaching $1$ on any nontrivial lattice hyperplane. See Bhattacharya–Rao [4, Theorem 20.1] and Kolassa [17, Chapter 4] for local lattice Edgeworth expansions under such conditions. ∎
A.2 Ratio of local CLT evaluations
We now deduce a uniform ratio expansion for probabilities at lattice points separated by a bounded shift $t$ with $\mathbf{1}^{\top}t=0$.
Lemma A.2 (Ratio lemma).
Assume the setting of Lemma A.1 and (34). Let $t\in\mathbb{Z}^{d}$ satisfy $\mathbf{1}^{\top}t=0$ and $\|t\|_{1}\leq 2$. Let $s\in\mathbb{Z}^{d}$ satisfy $\mathbf{1}^{\top}s=m$ and $\|(s-\mu)^{-}\|_{2}\leq m^{5/8}$. Then, for $m$ large enough (depending only on $d$ and $p$),
$\log\frac{\mathbb{P}(S_{m}=s-t)}{\mathbb{P}(S_{m}=s)}=\frac{1}{m}\,t^{\top}\Sigma_{\mathrm{ps}}^{+}(s-\mu)-\frac{1}{2m}\,t^{\top}\Sigma_{\mathrm{ps}}^{+}t+O(m^{-3/4}),$
where $\Sigma_{\mathrm{ps}}^{+}$ is the Moore–Penrose pseudoinverse of the per-sample covariance $\Sigma_{\mathrm{ps}}=(1/m)\mathrm{Cov}(S_{m})$ on $T$. In particular, in the fixed-composition setting one has $\Sigma_{\mathrm{ps}}=(1-\pi)\Sigma_{0}+\pi\Sigma_{1}$, consistent with (1) in the main text.
Moreover, if $\|(s-\mu)^{-}\|_{2}\leq M\sqrt{m}$ for a fixed constant $M$, then the same argument yields an improved remainder $O_{M}(1/m)$.
Proof.
Write $z^{-}=(s^{-}-\mu^{-})/\sqrt{m}$ and $z^{\prime-}=z^{-}-t^{-}/\sqrt{m}$. Under the assumed window, $\|z^{-}\|_{2}\leq m^{1/8}$ and $\|z^{\prime-}\|_{2}\leq 2m^{1/8}$ for large $m$, so Lemma A.1 applies at both $s$ and sts-t. Taking the ratio cancels the normalizing constant in (35) and yields
$\log\frac{\mathbb{P}(S_{m}=s-t)}{\mathbb{P}(S_{m}=s)}=-\frac{1}{2}\Bigl((z^{\prime-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}z^{\prime-}-(z^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}z^{-}\Bigr)+\Delta_{m},$
where $\Delta_{m}$ is the log-ratio of the Edgeworth correction factors.
For the Gaussian quadratic part, expand $z^{\prime-}=z^{-}-t^{-}/\sqrt{m}$ to obtain
$\frac{1}{2}\Bigl((z^{\prime-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}z^{\prime-}-(z^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}z^{-}\Bigr)=\frac{1}{m}(t^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}(s^{-}-\mu^{-})-\frac{1}{2m}(t^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}t^{-}.$
Since $t\in T$ and $s-\mu\in T$ (both have coordinate sum 0), bilinear forms in reduced coordinates agree with those formed using the pseudoinverse on $T$:
$(t^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}(s^{-}-\mu^{-})=t^{\top}\Sigma_{\mathrm{ps}}^{+}(s-\mu),\qquad(t^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}t^{-}=t^{\top}\Sigma_{\mathrm{ps}}^{+}t.$ (36)
To verify (36), note that the map $u\mapsto(u^{-},-\mathbf{1}^{\top}u^{-})$ is a linear isomorphism from $\mathbb{R}^{d-1}$ onto $T$, and $\Sigma_{\mathrm{ps}}$ restricted to $T$ is represented in these coordinates by $\Sigma_{\mathrm{ps}}^{-}$; hence solving $\Sigma_{\mathrm{ps}}x=y$ with $x,y\in T$ is equivalent to solving $\Sigma_{\mathrm{ps}}^{-}x^{-}=y^{-}$, which implies the stated identities.
It remains to bound $\Delta_{m}$. Write $E(z)=1+m^{-1/2}Q_{m}(z)+R_{m}(z)$ so that the Edgeworth factor in (35) is $E(z^{-})$. On the window $\|z^{-}\|_{2}\leq 2m^{1/8}$, Lemma A.1 gives $|m^{-1/2}Q_{m}(z)|\leq Cm^{-1/8}$ and $|R_{m}(z)|\leq Cm^{-1/4}$, hence $E(z)$ is bounded away from 0 and $\log E$ is Lipschitz on this range. Moreover, since $Q_{m}$ is cubic with uniformly bounded coefficients, its gradient is quadratic, so
$|Q_{m}(z^{\prime-})-Q_{m}(z^{-})|\leq C\|z^{-}\|_{2}^{2}\cdot\|z^{\prime-}-z^{-}\|_{2}\leq Cm^{1/4}\cdot m^{-1/2}=Cm^{-1/4}.$
Multiplying by $m^{-1/2}$ yields a contribution $O(m^{-3/4})$. The remainder difference $R_{m}(z^{\prime-})-R_{m}(z^{-})$ is smaller under the same window. Therefore $\Delta_{m}=O(m^{-3/4})$, uniformly over all admissible $s$ and $t$, and combining with the Gaussian part and (36) proves the claim. ∎
Lemma A.3 (Refined ratio lemma: isolating the Edgeworth-gradient term).
Assume the setting of Lemma A.1 and (34). Let $t\in\mathbb{Z}^{d}$ satisfy $\mathbf{1}^{\top}t=0$ and $\|t\|_{1}\leq 2$. Let $s\in\mathbb{Z}^{d}$ satisfy $\mathbf{1}^{\top}s=m$ and $\|(s-\mu)^{-}\|_{2}\leq m^{5/8}$, and write $z^{-}=(s^{-}-\mu^{-})/\sqrt{m}$. Then, for $m$ large enough (depending only on $d$ and $p$),
$\log\frac{\mathbb{P}(S_{m}=s-t)}{\mathbb{P}(S_{m}=s)}=\frac{1}{m}\,t^{\top}\Sigma_{\mathrm{ps}}^{+}(s-\mu)-\frac{1}{2m}\,t^{\top}\Sigma_{\mathrm{ps}}^{+}t+\frac{1}{m}\,\mathcal{Q}_{m,t}(z^{-})+\rho_{m,t}(s),$
where $\mathcal{Q}_{m,t}$ is a quadratic polynomial in $z^{-}$ with coefficients bounded in terms of $(d,p)$ and satisfying
$|\mathcal{Q}_{m,t}(z^{-})|\leq C\,(1+\|z^{-}\|_{2}^{2}),$
and the remainder satisfies
$|\rho_{m,t}(s)|\leq C\,m^{-3/2}\,(1+\|z^{-}\|_{2}^{9}).$
All constants depend only on $(d,p)$.
Proof.
We use a local lattice Edgeworth expansion to one order higher than Lemma A.1. Because the summands are uniformly bounded (categorical) and satisfy the strong aperiodicity condition implied by (34), local Edgeworth expansions of arbitrary order hold with polynomial-growth remainders; see Bhattacharya–Rao [4, Chapter 5] and Kolassa [17, Chapter 4]. In particular, there exist polynomials $Q^{(1)}_{m}$ (cubic) and $Q^{(2)}_{m}$ (degree at most $6$), with coefficients bounded in terms of $(d,p)$, such that uniformly over $\|z\|_{2}\leq 2m^{1/8}$,
$\mathbb{P}(S_{m}=s)=\frac{1}{(2\pi m)^{(d-1)/2}\sqrt{\det\Sigma_{\mathrm{ps}}^{-}}}\,\exp\!\left(-\frac{1}{2}(z^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}z^{-}\right)\,\Bigl(1+m^{-1/2}Q^{(1)}_{m}(z^{-})+m^{-1}Q^{(2)}_{m}(z^{-})+R^{(2)}_{m}(z^{-})\Bigr),$ (37)
where the remainder satisfies
$|R^{(2)}_{m}(z)|\leq C\,m^{-3/2}\,(1+\|z\|_{2}^{9}).$ (38)
(The exponent $9$ is not special; any sufficiently large fixed exponent works.)
Let $z^{-}=(s^{-}-\mu^{-})/\sqrt{m}$ and $z^{\prime-}=z^{-}-t^{-}/\sqrt{m}$. Apply (37) at both $s$ and sts-t and take the log-ratio. The Gaussian quadratic part yields
$\frac{1}{2}\Bigl((z^{\prime-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}z^{\prime-}-(z^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}z^{-}\Bigr)=\frac{1}{m}(t^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}(s^{-}-\mu^{-})-\frac{1}{2m}(t^{-})^{\top}(\Sigma_{\mathrm{ps}}^{-})^{-1}t^{-},$
which equals $\frac{1}{m}t^{\top}\Sigma_{\mathrm{ps}}^{+}(s-\mu)-\frac{1}{2m}t^{\top}\Sigma_{\mathrm{ps}}^{+}t$ by (36).
It remains to expand the ratio of the Edgeworth correction factors. Write
$E(z):=1+m^{-1/2}Q^{(1)}_{m}(z)+m^{-1}Q^{(2)}_{m}(z)+R^{(2)}_{m}(z),$
so that the correction factor in (37) is $E(z^{-})$. On the window $\|z\|_{2}\leq 2m^{1/8}$ we have $|m^{-1/2}Q^{(1)}_{m}(z)|\leq Cm^{-1/8}$, $|m^{-1}Q^{(2)}_{m}(z)|\leq Cm^{-1/4}$, and $|R^{(2)}_{m}(z)|\leq Cm^{-3/2}(1+\|z\|_{2}^{9})$, so $E(z)$ is bounded away from 0. Thus $\log E$ is Lipschitz and admits the Taylor expansion $\log(1+x)=x+O(x^{2})$ uniformly on this range.
Decompose
$\log\frac{E(z^{\prime-})}{E(z^{-})}=\underbrace{m^{-1/2}\bigl(Q^{(1)}_{m}(z^{\prime-})-Q^{(1)}_{m}(z^{-})\bigr)}_{(\star)}+\underbrace{m^{-1}\bigl(Q^{(2)}_{m}(z^{\prime-})-Q^{(2)}_{m}(z^{-})\bigr)}_{(\star\star)}+O\!\left(m^{-3/2}(1+\|z^{-}\|_{2}^{9})\right),$
where the last term collects the contribution of $R^{(2)}_{m}$ and the quadratic error $O(x^{2})$ from $\log(1+x)$. Since $Q^{(1)}_{m}$ is cubic with bounded coefficients, a first-order Taylor expansion gives
$Q^{(1)}_{m}(z^{\prime-})-Q^{(1)}_{m}(z^{-})=\nabla Q^{(1)}_{m}(z^{-})\cdot(z^{\prime-}-z^{-})+O\!\left(\|z^{\prime-}-z^{-}\|_{2}^{2}(1+\|z^{-}\|_{2})\right),$
and $z^{\prime-}-z^{-}=-t^{-}/\sqrt{m}$. Therefore
$(\star)=-\frac{1}{m}(t^{-})^{\top}\nabla Q^{(1)}_{m}(z^{-})+O\!\left(m^{-3/2}(1+\|z^{-}\|_{2})\right).$
Define the quadratic polynomial $\mathcal{Q}_{m,t}(z):=-(t^{-})^{\top}\nabla Q^{(1)}_{m}(z)$. Because $\nabla Q^{(1)}_{m}$ is quadratic with bounded coefficients, $|\mathcal{Q}_{m,t}(z)|\leq C(1+\|z\|_{2}^{2})$.
For $(\star\star)$, since $Q^{(2)}_{m}$ has degree at most $6$ with bounded coefficients, a first-order Taylor bound yields
$|Q^{(2)}_{m}(z^{\prime-})-Q^{(2)}_{m}(z^{-})|\leq C(1+\|z^{-}\|_{2}^{5})\|z^{\prime-}-z^{-}\|_{2}\leq C(1+\|z^{-}\|_{2}^{5})m^{-1/2},$
so $(\star\star)=O(m^{-3/2}(1+\|z^{-}\|_{2}^{5}))$. Combining these estimates gives
$\log\frac{E(z^{\prime-})}{E(z^{-})}=\frac{1}{m}\mathcal{Q}_{m,t}(z^{-})+O\!\left(m^{-3/2}(1+\|z^{-}\|_{2}^{9})\right).$
Adding the Gaussian part completes the proof. ∎